### v1

# Deterministic Dynamic Edge-Colouring

Abstract
Given a dynamic graph $G$ with $n$ vertices and $m$ edges subject to insertion an deletions of edges, we show how to maintain a $(1+\varepsilon)\Delta$-edge-colouring of $G$ without the use of randomisation. More specifically, we show a deterministic dynamic algorithm with an amortised update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ using $(1+\varepsilon)\Delta$ colours. If $\varepsilon^{-1}\in 2^{O(\log^{0.49}n)}$, then our update time is sub-polynomial in $n$.
While there exists randomised algorithms maintaining colourings with the same number of colours [Christiansen STOC’23, Duan, He, Zhang SODA’19, Bhattacarya, Costa, Panski, Solomon SODA’24] in polylogarithmic and even constant update time, this is the first deterministic algorithm to go below the greedy threshold of $12\Delta-$ 1 colours for all input graphs. On the way to our main result, we show how to dynamically maintain a shallow hierarchy of degree-splitters with both recourse and update time in $n^{o(1)}$. We believe that this algorithm might be of independent interest.
1 Introduction & Related Work
Edge-colouring is a classic and fundamental problem studied extensively in both graph theory as well as many different models of computation [${}^{+}$85, Ber22, Chr23, DHZ19, BGW21]. Given a simple graph $G=(V,E)$, a $k$-edge-colouring of $G$ is a function $c:E(G)\rightarrow[k]$ that maps every edge to a colour in some colour set $[k]=\{1,2,\dots,k\}$, An edge-colouring is proper if $e\cap e^{\prime}\neq\emptyset$ implies that $c(e)\neq c(e^{\prime})$ for any pair of distinct edges $e,e^{\prime}$. The chromatic index $\chi^{\prime}(G)$ of $G$ is then the smallest $k$ for which a proper $k$-edge-colouring of $G$ exists. It is not difficult to check that $\chi^{\prime}(G)\geq\Delta$ if $G$ has maximum degree $\Delta$. A classic result of Vizing [Viz64] shows that in fact $\chi^{\prime}(G)\leq\Delta+1$. Deciding whether $\chi^{\prime}(G)=\Delta$ or $\chi^{\prime}(G)=\Delta+1$ is, however, not easy, and it was shown by Holyer that determining $\chi^{\prime}(G)$ exactly is, in fact, NP-complete [Hol81].
Recently, the edge-colouring problem has received a lot of attention in the dynamic algorithms community. Here one studies dynamic graphs, where edges are continually inserted and deleted, and the challenge is to maintain some proper edge-colouring of the graph as it is subjected to updates. There has been some exciting progress on developing efficient randomised dynamic algorithms using only $(1+\varepsilon)\Delta$ colours. This number of colours go significantly below the greedy-threshold of $12\Delta-$ 1 colours. Duan, He, and Zhange [DHZ19] showed how to maintain such a $(1+\varepsilon)\Delta$-edge-colouring with an update time of $O(\tfrac{\log^{7}n}{\varepsilon^{2}})$ provided that $\Delta=\Omega(\tfrac{\log^{2}n}{\varepsilon^{2}})$. Christiansen got rid of all requirements on $\Delta$ while maintaining a worst-case update time in $O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$. Very recently Bhattacarya, Costa, Panski, and Solomon [BCPS] pushed the update time all the way down to $\tilde{O}(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1}))$ under some additional assumptions on $\Delta$ and $\varepsilon$.
However, all of the above algorithms rely on randomness in several and crucial ways, and the landscape of deterministic and dynamic algorithms is a lot less colourful. Barenboim and Maimon [BM17] gave a dynamic algorithm with $O(\sqrt{\Delta})$ worst-case update time using $O(\Delta)$ colours. This was subsequently improved by Bhattacarya, Chakrabarty, Henzinger and Nanongkai [BCHN18] who gave a fully dynamic algorithm for computing a $(2\Delta-1)$-edge-colouring with worst case $O(\log\Delta)$ update time. Bhattacarya, Costa, Panski, and Solomon [BCPS23a] and independently Christiansen, Rotenberg, and Vlieghe [CRV23] showed how to maintain a $(\Delta+O(\alpha))$-edge-colouring in amortised $O(\operatorname{\text{{\rm poly}}}(\log n))$ update-time, where $\alpha$ is the arboricity of $G$. However, if $\Delta\sim\alpha$, this algorithm will not always use a number of colours below the greedy threshold as $\alpha$ can be as big as $\tfrac{\Delta}{2}$. In particular, until now no deterministic and dynamic algorithm always using less than $12\Delta-$ 1 colours with sub-linear update time was known.
In general it is an interesting and challenging task to design deterministic algorithms with guarantees similar to the randomised ones. In this paper, we will address this gap between randomised and deterministic dynamic algorithm for $(1+\varepsilon)\Delta$-edge-colouring by providing a deterministic and fully-dynamic algorithm for maintaining a $(1+\varepsilon)\Delta$-edge-colouring in $n^{o(1)}$ amortised update-time:
Theorem 1 (Informal).
Let $G$ be a dynamic graph and $\varepsilon^{-1}\leq 2^{\operatorname{\text{{\rm poly}}}(\log\log n)}$ a given parameter. Then, we can maintain a proper $((1+\varepsilon)\Delta)$-edge-colouring of $G$ in $n^{o(1)}$ amortised update time per operation.
1.1 Main technical challenges
In this section, we will pinpoint two issues towards designing deterministic dynamic edge-colouring algorithms with few colours. These are barriers we know how to overcome in either the static setting or by using randomisation, but so far, to the best of our knowledge, no current techniques deal with these issues directly.
Many static algorithms, distributed algorithms, and dynamic randomised algorithms produce edge-colourings with few colours via a two-step approach. The first step is a form for dimenstionality reduction; one attempts to reduce the problem to a case where $\Delta$ is small. The second step is then to apply an algorithm that is fast and efficient for graphs with small maximum degree.
Many static algorithms [${}^{+}$85, BCPS23b, Sin19] and many distributed algorithms [GKMU18, ${}^{+}$20, GS17] adress the first step by computing a hierarchy of degree-splitters. A degree-splitter splits a graph $G$ into two covering subgraphs $G_{1},G_{2}$ such that the maximum degree of $G_{i}$ is roughly half of the maximum degree of $G$. By applying this procedure recursively one can reduce the problem to $\tfrac{\Delta}{\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)}$ sub-problems each of maximum-degree $\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)$. However, these static and precise approaches for splitting the degree, seems less compatible with changes to the graph.
If the maximum degree is large enough, it is easy to realise something similar in the randomised setting, even dynamically against an oblivious adversary. Indeed, one can place each edge into one of $\tfrac{\Delta}{\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)}$ subgraphs uniformly at random. It then follows from a standard application of Chernoff-bounds that the maximum degree of each such subgraph is split nicely. A different approach for reducing $\Delta$ is to sample uniformly at random a palette $S\subset[t]$ of colours such that the graph induced by all edges coloured with a colour from $S$, say $G[S]$, exhibits useful properties. Indeed, one can show that with high-probability the maximum degree of $G[S]$ is small and every vertex in $G$ has a free colour in $S$. One can then colour a new edge, by viewing it as being inserted into $G[S]$. This is the approach taken in [Chr23, DHZ19]. It is, however, not clear how to perform such a reduction in the deterministic dynamic setting; naively building a hierarchy can easily result in an insurmountable recourse (see section 2), and sampling colours or edges relies crucially on the power of randomisation.
The second challenge is to efficiently colour a low-degree graph. In the static setting one could rely on parallelism [${}^{+}$85], and in the dynamic & distributed settings recent progress has shown that the recourse for extending colourings on low-degree graphs is small [Ber22, BD23, Chr23, DHZ19] and, more importantly, can be efficiently realised using randomisation [Ber22, BD23, Chr23, DHZ19]. However, naively searching for such an extending sequence of recolourings can be expensive, and so it is unclear how to locate it without using randomisation.
1.2 Further Related Work
Gabow, Nishizeki, Kariv, Leven, and Tereda [${}^{+}$85] showed how to compute a $(\Delta+1)$-edge-colouring in $\tilde{O}(m\sqrt{n})$ or $\tilde{O}(m\Delta)$ time. This improved over the $O(mn)$ time algorithm implied by the work of Vizing [Viz64]. Recently this was improved by Sinnamon [Sin19], and in special settings by Bernshteyn and Dhawan [BD23] and by Bhattacarya, Costa, Panski, and Solomon [BCPS23b]. Alon [Alo03] and Cole, Ost and Schirra [COS01] gave a near linear-time algorithm for $\Delta$-edge-colouring bipartite graphs. By applying a reduction due to Karloff and Shmoys [KS87], these algorithms yield $3\lceil\frac{\Delta}{2}\rceil$-edge-colouring algorithms for general graphs. When $\Delta=\Omega{(\varepsilon^{-1}\log n)}$, Duan, He and Zhang [DHZ19] gave a $O(\varepsilon^{-2}m\log^{6}n)$ time algorithm for computing a $(1+\varepsilon)\Delta$-edge-colouring. This has recently been improved by Bhattacarya, Costa, Panski, and Solomon [BCPS] to $\tilde{O}(m\operatorname{\text{{\rm poly}}}(\varepsilon^{-1}))$ time for many values of $\Delta$ and $\varepsilon$.
The edge-colouring problem has also been extensively studied in the LOCAL model [Ber22, BD23, ${}^{+}$18, Chr23, GKMU18, ${}^{+}$20, GS17, SV19] (see for instance [${}^{+}$18, GKMU18] for extensive surveys). Some of these algorithms rely on degree-splitting either rely on or directly consider the degree-splitting problem [GKMU18, ${}^{+}$20, GS17].
1.3 Results
Our main result it the following theorem:
Theorem 2.
Let $G$ be a dynamic graph, and $10<\varepsilon<$ 1 a given parameter. We can maintain a proper $(1+\varepsilon)\Delta$-edge-colouring of $G$ in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ amortised [^1] Here $\tilde{O}_{\log\varepsilon^{-1}}$ hides $\log\log n$ and $\log\varepsilon^{-1}$ factors. update time per operation.
On the way to this result, we present several algorithms which we believe might be of independent interest. First of all, we design what we call a dynamic $t$-splitter; a dynamic algorithm maintaining a $t$-edge-colouring such that the graph induced by each colour has low-degree.
Theorem 3.
Let $G$ be a dynamic graph, and suppose $\varepsilon>0$, $t\in\mathbb{N}$, and an upper bound $\Delta_{\max}$ on the maximum degree of $G$ through the entire update sequence is given. Assume that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then there is a dynamic algorithm that computes a $t$-splitter of $G$ with an amortised update-time of $O(t^{3}\varepsilon^{-4}\log^{3}m)$.
Second of all, we show how to build a shallow hierarchy of $t$-splitters that splits the degree all the way down to $\operatorname{\text{{\rm poly}}}(\log n,\varepsilon^{-1})$. We believe that this might be a useful algorithmic tool for developing efficient dynamic algorithms for other problems as well.
Theorem 4.
Let $G$ be a dynamic graph, and let $01>\varepsilon>$ 0 and an upper bound $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ on the maximum degree of $G$ through the entire update sequence be given. Set $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$ and $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Then there exists a dynamic algorithm maintaining parameters $h,i,j$ with $h=i+j\leq 4\sqrt{\log n}$ and a set of graphs $\mathcal{G}_{h}$ satisfying that:
  1. 1.
    $|\mathcal{G}_{i+j}|\leq t_{1}^{i}t_{2}^{j}$
  2. 2.
    $\hat{\Delta}_{i+j}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
The algorithm has amortised recourse and update time both in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$.
Finally, we show how to efficiently $(\Delta_{\max}+1)$-edge-colour graphs with low maximum degree:
Theorem 5.
Let $G$ be a dynamic graph, and let $\Delta_{\max}$ be a known upper-bound on the maximum degree throughout the entire update sequence. Then, we can maintain a proper $(\Delta_{\max}+1)$-edge-colouring of $G$ in $(\Delta_{\max}+1)^{\tilde{O}(\sqrt{\log n})}$ worst-case update time per operation.
2 Informal Overview
In this section, we will give a high-level overview of the main ideas and the algorithm. We first introduce the overall approach:
2.1 The overall approach
From now on, unless otherwise noted, we will assume that we have a known upper bound $\Delta_{\max}$ on the maximum degree $\Delta$. This assumption will later be removed by using a degree-scheduler identical to the one used by Duan, He, and Zhang [DHZ19].
We will take a two-step approach that, at a high-level, is inspired by the distributed algorithm of Ghaffari, Kuhn, Maus, and Uitto [GKMU18]. The idea is to partition the graph into a set of at most $\tfrac{\Delta}{t}$ subgraphs $\mathcal{G}_{h}$ such that $\{E(G_{i})\}_{G_{i}\in\mathcal{G}_{h}}$ partitions $E(G)$ and such that $\Delta(G_{i})\leq(1+\tfrac{\varepsilon}{2})\cdot{}t$ for all $G_{i}\in\mathcal{G}_{h}$. Then one can colour each $G_{i}\in\mathcal{G}_{h}$ with a separate palette using some algorithm that is efficient in low-degree graphs, and then combine all of these colourings into a $(1+\varepsilon)\Delta$-edge-colouring of the original graph $G$.
As a warm-up, we briefly sketch how to boot-strap a simple static algorithm for computing a $(1+\varepsilon)\Delta$-edge-colouring in near-linear time using the algorithms of Gabow et Al. [${}^{+}$85]. This algorithm was recently, and independently, obtained by Elkin and Khuzman [EK24], while this manuscript was under preparation, and we therefore refer to their manuscript for a formal description of the algorithm.
The idea is the following: a degree-splitter with discrepancy $\delta$ is a partition of the edge set into two sets $E_{1}$ and $E_{2}$, such that for all vertices $v$, we have that $d_{G[E_{1}]}(v),d_{G[E_{w}]}(v)\leq\tfrac{\Delta}{2}+\delta$, where $G[E_{i}]$ is the graph induced in $G$ by the set $E_{i}$ i.e. the subgraph of $G$ containing all edges of $E_{i}$. This is a relaxed definition of degree-splitting similar to the one used in [GS17], but other variants have also been studied (see for example [${}^{+}$20]). Given an efficient degree-splitter with small enough discrepancy, one can easily compute a partition $\mathcal{G}_{h}$ of the desired form by recursively applying the degree splitter to $G_{1}$ and $G_{2}$ to build a hierarchy whose lowest level form $\mathcal{G}_{h}$.
Gabow et Al. [${}^{+}$85] showed how to compute a degree-splitter with a discrepancy of only $2$ in the static setting. By recursively applying this degree splitter roughly $h=\log\Delta-\log\tfrac{\log n}{\varepsilon}$ times (or 0 times if $\Delta=O(\tfrac{\log n}{\varepsilon})$) one can achieve a split into roughly $\tfrac{\Delta}{\tfrac{\log n}{\varepsilon}}$ subgraphs, each with maximum degree at most $\tfrac{\log n}{\varepsilon}+O(\log n+\sum_{i=0}^{\infty}\tfrac{2}{2^{i}})$. By rescaling $\varepsilon$ a bit, this yields the desired partition. Finally, one can colour each graph in $\mathcal{G}_{h}$ with a distinct palette efficiently using another algorithm of Gabow et Al. [${}^{+}$85].
As we have already briefly touched upon in section 1.1, the two main challenges with making such an approach dynamic are the following: 1)1)1 ) how to maintain the degree splitters, and 2)2)2 ) how to efficiently colour the low-degree graphs.
As for 1)1)1 ), we note that first of all, as far as we are aware, no deterministic dynamic degree-splitting algorithm is known. While it is easy to devise a randomised such algorithm against an oblivious adversary for graphs with not too small maximum degree (simply partition the edges u.a.r into $2$-edge sets), it is unclear how to do this deterministically. Second of all, if a deterministic degree-splitter has recourse at least $10$, then naively maintaining a hierarchy of degree-splitters becomes infeasible as the recourse at depth $h$ of the hierarchy can be as large as $10^{h}$. In order to split the degree from $n^{0.75}$ to $n^{0.25}$, we would require $h\geq\tfrac{1}{2}\log n$, thus resulting in super-linear recourse, and therefore a super-linear update time.
We overcome this as follows: instead of designing a dynamic degree-splitter, we consider the dynamic $t$-splitting problem, where we are interested in splitting the graph into $t$-subgraphs, each with some restrictions on the maximum degree:
Problem 6.
Given a dynamic graph $G$ subject to insertions and deletions, maintain a colouring $c:E(G)\rightarrow[t]$ such that for all vertices $v\in V(G)$ and all colours $i\in[t]$, we have that the number of edges coloured $i$ incident to $v$, $d_{i}(v)$, is upper bounded by $d_{i}(v)\leq(1+\varepsilon)\frac{\Delta}{t}$.
We show in Section 4 that we can, in fact, maintain a dynamic $t$-splitter in $\tilde{O}(\operatorname{\text{{\rm poly}}}(t,\log(n),\varepsilon))$ update time, but crucially with only $\operatorname{\text{{\rm poly}}}(\log(n),\varepsilon)$ recourse. This data-structure is far from trivial, and requires many ideas of its own. We will discuss it in Section 2.2.
Since the $t$-splitter has low recourse, we can tune $t$ very aggressively. In turn, we can get a dynamic $\frac{\Delta}{\operatorname{\text{{\rm poly}}}(\log n,\varepsilon^{-1})}$-splitter by maintaining a very shallow hierarchy of correctly tuned $t$-splitters. Indeed, we can set $t\sim 2^{\sqrt{\log n}}$ and split the degree down to $\sim 2^{\sqrt{\log n}}$ using a hierarchy of depth only $h\sim\sqrt{\log n}$. This allows us to bound the recourse $R(h)$ at depth $h$ as:
$R(h)\leq\operatorname{\text{{\rm poly}}}(\log(n),\varepsilon)^{h}=2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$
Then, we can set $t^{\prime}\sim\operatorname{\text{{\rm poly}}}(\log n,\varepsilon^{-1})$ to further split the degree down to $\operatorname{\text{{\rm poly}}}(\log n,\varepsilon^{-1})$, while only increasing the recursive depth by an extra additive factor of $\sqrt{\log n}$. Since each update takes only $\tilde{O}(\operatorname{\text{{\rm poly}}}(t,\log(n),\varepsilon))$ time, this allows us to split the degree down to $\operatorname{\text{{\rm poly}}}(\log n,\varepsilon^{-1})$ with both recourse and update time in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$.
In order to address 2)2)2 ), we will make use of multi-Step Vizing chains, a term coined by Bernshteyn [Ber22] and a topic that has seen a lot of recent development [Ber22, BD23, GP20, Chr23]. Here, one attempts to find a sequence of recolourings that creates an uncoloured edge with a free colour. The idea is to construct such a sequence of recolourings by first constructing a $1$-step Vizing chain, which is a construction that goes back all the way to Vizing’s original paper [Viz64]: a Vizing chain consists of two parts: a) a fan $F$ of edges and b) a bichromatic path $P$ of edges coloured with two colours $\kappa_{1}$ and $\kappa_{2}$. The fan $F$ consists of a center vertex $u$ together with edges $uw_{1},\dots,uw_{k}$ such that the colour of $uw_{i+1}$ is available at $w_{i}$, meaning that no edges incident to $w_{i}$ has the colour $c(uw_{i+1})$. The bichromatic path $P$ then has to begin at $w_{k}$ and consist only of edges coloured $\kappa_{1}$ and $\kappa_{2}$, where $\kappa_{1}$ is available at $u$ and $\kappa_{2}$ is available at $w_{k}$. A Vizing chain can be shifted by recolouring edges in $F$ and subsequently in $P$ as shown in Figure 1 to the left.
Figure 1: An augmenting Vizing chain is shifted on the left, and a truncated Vizing chain is shifted on the right. This illustration is from [Chr23]
Observe that if $P$ is not too long, we are already happy with the construction, as we need only recolour few edges. However, $P$ could potentially have length $\Omega(n)$. In this case, we can truncate it by uncolouring an edge along $P$ and only shifting the first part of $P$ as shown in Figure 1 to the right.
Then we may build a new Vizing chain on top of the new uncoloured edge in the hope that this chain has a shorter length. We may iterate this to construct multi-step Vizing chains (see Figure 2), as has been done in similar fashion in [Ber22, BD23, Chr23, GP20]. The goal is to get a short Vizing chain which is augmenting in the sense that after shifting it, the remaining uncoloured edge has a free colour. In order to more readily reason about these chains, we will require them to be non-overlapping in the sense that we impose restrictions on how the $i^{\text{th}}$ extension can overlap with earlier chains.
Figure 2: On the left, we see a non-overlapping multi-step Vizing chain. On the right, we see a multi-step Vizing chain overlapping in a single edge. This illustration is from [Chr23]
The main issue is that all known results about Vizing chains are either existential or obtained via randomised constructions. As noted by Bernshteyn & Dhawan [BD23] naively applying the argument of Christiansen [Chr23] to search for a Vizing chain can result in linear update time, and so it seems hopeless to perform some deterministic search for a short and augmenting Vizing chain. However, we will show that one can efficiently maintain data structures contianing enough information to be able to locate certain classes of short and augmenting Vizing chains efficiently. The main observation is that, under the correct tuning of parameters, one has that while one uncoloured edge can be part of many different multi-step Vizing chains, not too many vertices can actually reach the edge via a multi-step Vizing chain. This means that assigning an edge a new colour, does not result in the need of updating too many different data-structures. While maintaining these data structures is intuitively quite simple, it is also quite technical. Therefore, we will give a separate overview of how to maintain and use these data structures in Section 2.3.
2.2 Dynamic $t$-splitters
In this section, we will give an overview of how to maintain a dynamic $t$-splitter with low recourse. We begin by sketching an algorithm for the case $t=2$. We note already now that this case is considerable easier than the case $t>2$. For the case $t=2$ , the algorithm has worst-case guarantees, but for $t>2$ all guarantees are amortised. After sketching the approach, we discuss the technical difficulties with generalising this approach, and sketch how we overcome them.
We will colour the graph with $2$ colours, colour $1$ and colour $2$. For a vertex $v$, we define its surplus of colour $i$ to be $s_{i}(v)=\max\{d_{i}(v)-\frac{\Delta_{\max}}{t},0\}$. We will maintain a colouring satisfying the following local invariant for some parameter $\eta$:
Invariant 7.
For all $uv\in E(G)$, we have that $c(uv)=i$ implies and all $i\in[t]$, we have that
$\max\{s_{i}(u),s_{i}(v)\}\leq\eta+\max\{s_{i+1}(u),s_{i+1}(v)\}$
where addition is performed modulo $2$.
This invariant is useful for two reasons: first of all, one can show that if it holds everywhere, then the maximum surplus is never too big, and second of all, whenever an edge $uv$ is inserted one can restore the invariant by only recolouring edges along a path with alternating colours. These paths bear some resemblance to the augmenting paths used in a distributed algorithm due to Ghaffari and Su [GS17].
Indeed, to see that the invariant implies a low maximal surplus, observe that if $s_{i}(v)$ is large, then any vertex $w$ neighbouring $v$ via an edge coloured $i$ must have large surplus $s_{i+1}(w)\geq s_{i}(v)-\eta$. Applying this approach recursively yields set $V_{1},V_{2},\dots,V_{k}$ where $u\in V_{j}$ implies that $s_{i+j-1\mod 2}(u)\geq s_{i}(v)-(j-1)\eta$. In particular, one has:
$|E(G[\cup_{h=1}^{k}V_{h}])]\geq|\cup_{h=1}^{k}V_{h}|\cdot{}(s_{i}(v)-(k-1)\eta)$
Indeed, observe that one double counts no edges as vertices in $V_{j}$ cannot be connected by an edge coloured $i+j-1\mod 2$, without violating the local invariant.
For a reasonable choice of $k,\eta$ and $s_{i}(v)$, we have that $s_{i}(v)-(k-1)\eta>(1+\tfrac{\varepsilon}{16})\tfrac{\Delta}{2}$. In particular, since the maximum subgraph density $\rho(G)=\max_{J\subset G}\{\frac{|E(J)|}{|V(J)|}\}\leq\frac{\Delta}{2}$ by the hand-shaking lemma, it follows by induction that $|\cup_{h=1}^{k}V_{h}|\geq(1+\tfrac{\varepsilon}{16})|\cup_{h=1}^{k-1}V_{h}|$. Hence, we conclude that $k=O(\tfrac{\log n}{\varepsilon})$, which naturally bounds how big $s_{i}(v)$ can be.
In order to see that one need only recolour edges along an alternating path, assume the edge $uv$ is inserted and receives colour $i$ minimizing $\max\{s_{i}(u),s_{i}(v)\}$. The choice of $i$ readily implies that at most one of $s_{i}(u)$ and $s_{i}(v)$ can be larger than $\frac{\Delta_{\max}}{2}$. If $s_{i}(w)\leq\frac{\Delta_{\max}}{2}$ after $uv$ is coloured by $i$, then no edge incident to $w$ will suddenly violate the invariant. So we need only focus on edges incident to the endpoint with the larger surplus in $i$. WLOG say $s_{i}(u)>\frac{\Delta_{\max}}{2}$. If so, then only edges coloured $i$ incident to $u$ can become violating. If an edge $ux$ becomes violating, then it must be that
$\max\{s_{i}(u),s_{i}(x)\}>\eta+\max\{s_{i+1}(u),s_{i+1}(x)\}$
Hence, recolouring $ux$ to colour $i+1\mod 2$ restores the invariant at all edges incident to $u$, since $s_{i+1}(u)<\frac{\Delta_{\max}}{2}$. Now some edges coloured $i+1$ incident to $x$ might be violating. These can be fixed as above. Note that since $s_{i+1}(x)\leq s_{i}(u)-\eta$, this process cannot be iterated too many times before one visits a vertex with less than $\eta$ surplus, and one restores the invariant. In order to implement the above efficiently, one can locate violating in-edges by updating information lazily via a round-robin scheme. The use of the round-robin scheme is inspired by a sequence of recent work on dynamic algorithms for maintaining an estimate of the maximum subgraph density $\rho$, see [SW20, ${}^{+}$], where a lazy round robin scheme is used to locate a chain of directed edges to re-orient.
In order to generalise the above to $t>2$, we will work with the following local invariant instead:
Invariant 8.
For all $uv\in E(G)$ and all $i\in[t]$, we have that
$L(uv)=s_{c(uv)}(u)+s_{c(uv)}(v)\leq\eta+s_{i}(u)+s_{i}(v)$
Many of the difficulties that arise when generalising the above approach to $t>2$ stem from the fact that the problem becomes a lot less symmetric. The generalisation of the local invariant is chosen carefully to account for this asymmetry. Again we wish to show that 1)1)1 ) the local invariant implies that the colouring is $t$-splitter, and 2)2)2 ) that we can restore the invariant in some efficient manner.
As for 1)1)1 ) the asymmetry is reflected in the fact that one can no longer extract dense subgraphs as in the argument above. Indeed, for $t>2$ it is unclear how to extend a density based argument. Before there was a natural way of assigning a high-surplus edge to the corresponding high-surplus endpoint. This allowed us to extract a dense subgraph, which we could use to show that the maximum surplus could not be too large. However, for $t>2$ it is unclear which endpoint one should assign a high-surplus edge to; indeed assigning it to the high-surplus vertex as before risks that some vertices do not get assigned more than $\tfrac{\Delta_{\max}}{2}$ edges. Furthermore, there is the additional caveat that it might not even be well-defined which endpoint has high-surplus, as both endpoints could contribute equally to $s_{c(uv)}(u)+s_{c(uv)}(v)$.
Instead, we show how in Section 4 that if $s_{c(uv)}(u)+s_{c(uv)}(v)$ is large, then one can extract a sequence of bipartite subgraphs $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}$, possibly with low density, containing an exponentially increasing number of vertices. This again allows us to bound $s_{c(uv)}(u)+s_{c(uv)}(v)$, and show that we indeed maintain a $t$-splitter.
The asymmetry is also reflected in 2)2)2 ). Before one could restore the invariant by recolouring a path, since fixing violating edges coloured $i$ did not cause edges coloured $i+1$ at the same vertex to become violating. However, if one has multiple colours to choose from, one might have to recolour multiple edges incident to each vertex. Hence, one has to analyse the recourse completely differently. We show that each recolouring of a violating edge cause the following potential to drop:
$\Phi(G,c)=\sum\limits_{i=1}^{t}\sum\limits_{v\in V(G)}\sum\limits_{j=1}^{s_{i}(v)}j$
thus allowing us to bound the recourse of the algorithm in an amortised fashion.
Finally, we arrive at the dynamic $t$-splitter by carefully adjusting the lazy round-robin schemes and the data structures used to account for the asymmetry and the fact that $t>2$.
2.3 $(\Delta_{\max}+1)$-edge-colouring of low degree graphs
As a warm-up we will informally sketch how to maintain a $(\Delta_{\max}+1)$-edge-colouring with $\tilde{O}(\operatorname{\text{{\rm poly}}}(\Delta_{\max})\cdot{}n^{\tfrac{2}{3}})$ update-time. Then we sketch how to generalise this approach in order to achieve the faster update time of $(\Delta_{\max}+1)^{o(1)}$.
The starting point is this: suppose that given any proper partial $(\Delta_{\max}+1)$-edge-colouring and an uncoloured edge $e$, there exists an augmenting, non-overlapping $3$-step Vizing chain with length $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})\cdot{}n^{\tfrac{1}{3}})$. The existence of such chains is easy to derive from the work of Christiansen [Chr23]. Technically speaking, these derivations used $\Delta$ and not $\Delta_{\max}$, but it is straight-forward to check that one can replace $\Delta$ by the less restrictive $\Delta_{\max}$. Since the final algorithm works with more steps, we do not explicitly derive this bound, but it follows from near-identical arguments to the proof of Corollary 46.
While a naive search for such a chain might take $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n)$ time, one can readily observe that any single edge can only be part of a $3$-step Vizing chains beginning at one of $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ different points. Indeed, any coloured edge is part of at most $\Delta_{\max}+1$ different bichromatic paths and at most 2 different fans. Since any bichromatic path can be a part of at most $2(\Delta_{\max}+1)$ different 1-step Vizing chains, the statement follows for $1$-step Vizing chains. To see that the statement holds for $3$-step Vizing chains, observe that in order to extend the chain through $e$, one has to be able to reach an edge incident to an edge sharing an endpoint with $e$. Since there are at most $2(\Delta_{\max}+1)$ such edges, one can apply the above argument again to see that there are at most $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ choices for the starting point of $2$-step Vizing chains. Iterating this argument then gives a similar bound for $3$-step Vizing chains.
The above argument relies crucially on the fact that we require the Vizing chain to be non-overlapping (see Definition 11 for a formal definition). Indeed, an issue could be that shifting the first step of a Vizing chain could create many new bichromatic paths not accounted for by the above argument. But for that to happen, the extension must circle back and overlap or intersect an ealier Vizing chain. This in turn means that the Vizing chain becomes overlapping, and thus is not considered for the above argument. This type of argument is not new, and has been used for instance by Bernshteyn [Ber22] and Christiansen [Chr23] to analyse the recourse of certain multi-step Vizing chain constructions.
In this paper, we use this fact very explicitly in our algorithm. The idea is to explicitly maintain enough information about bichromatic paths in the graph to be able to construct an augmenting and non-overlapping $3$-step Vizing chain fast: if each vertex knows the $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n^{\tfrac{1}{3}})$ shortest augmenting $1$-step Vizing chains going through it, then one can update information about the shortest augmenting and non-overlapping $2$-step Vizing chain fast: simply consider all $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n^{\tfrac{1}{3}})$ different extension points, locate the shortest augmenting $1$-step Vizing chains and extend via the shortest such chains. We can then update information about $3$-step Vizing chains in a similar fashion.
Since any edge is part of only $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ colouring, re-colouring, or un-colouring such an edge only affects at most $O(\operatorname{\text{{\rm poly}}}(\Delta_{\max}))$ data-structures. By determining all data-structures containing augmenting $1$-step Vizing chains and augmenting $2$-step Vizing chains, we can first update all of the $1$-step structures and then use them to update the $2$-step structures efficiently, and so on.
The above approach has some glaring issues: first of all, we need the extensions to be non-overlapping. This requires us to check whether each extension is overlapping. However, by extending in a controlled manner [Ber22, GP20], one can ensure that not too many extension are overlapping, and thus one can actually afford to check for each chain whether it is overlapping, in $\tilde{O}(\operatorname{\text{{\rm poly}}}(\Delta_{\max})n^{\tfrac{1}{3}})$ time per chain. By carefully choosing how many augmenting chains need to be saved for each number of steps, one can ensure that one actually finds the shortest such chains.
Secondly, changing the colour of edges might cause some of the fan-constructions to become out-dated, and so explicitly storing the Vizing chains might suddenly require information about a different type of extension. To remedy this issue, we consider a stepping process, which emulates the structure of consistent Vizing chains: in each step the process is allowed to jump to a neighbouring vertex and then follow a new bichromatic path. This allows us to maintain information about processes for many different kind of extensions, which in turn allows us to quickly update the data structures whenever a fan-construction becomes out-dated.
The final algorithm is then similar to above, but instead we will work with step-length of roughly $2^{O(\sqrt{(}\log n))}$ and $O(\sqrt{\log n})$ steps. By carefully designing data-structures, we show that then the above approach can in fact be generalised to yield a dynamic algorithm with $(\Delta_{\max}+1)^{\tilde{O}(\sqrt{(}\log n))}$ worst-case update time.
2.4 Outline of the paper
In Section 3, we define notation and discuss some preliminaries. Then in Section 4, we present our dynamic $t$-splitter, before we show how to build a hierarchy of such splitters in Section 5. In Section 6, we present our algorithm for colouring low-degree graphs. Finally, in Section 7 we combine all of the above to get our final algorithm.
3 Preliminaries & Notation
The first parts of this section is very similar to Section 2 of the full-version of [Chr23], as we need similar terminology. For a positive integer $t$, we let $[t]=\{1,2,\dots,t\}$ be the set of numbers in $\mathbb{N}$ that are greater than 00, but at most $t$. We let $G=(V,E)$ be a graph on $n$ vertices and $m$ edges. A subgraph $H\subset G$ of $G$ is a graph such that $V(H)\subset V(G)$ and $E(H)\subset E(G)$. The $t$-hop neighbourhood of $H$ in $G$, $N^{t}(H)$ consists of all of the vertices in $G$ of distance at most $t$ (in $G$) to a vertex in $H$. In particular, $H\subset N(H)$. A proper (partial) $k$-edge-colouring $c$ of $G$ is a function $c:E(G)\mapsto[k]\cup\{\neg\}$ such that if $c(e)\neq\neg$ and $c(e^{\prime})\neq\neg$, then $e\cap e^{\prime}$ implies that $c(e)\neq c(e^{\prime})$. An edge $e$ such that $c(e)=\neg$ is said to be uncoloured. Two uncoloured edges may share an endpoint in a proper partial colouring. If no edges are uncoloured, we say that the colouring is a proper $k$-edge-colouring. For a vertex set $U\subset V(G)$ (resp. edge set $U^{\prime}\subset E(G)$), we let $G[U]$ (resp. $G[U^{\prime}]$) be the induced graph of $U$ in $G$ i.e. the subgraph of $G$ on vertex set $V(G[U])=U$ containing the maximum number of edges (resp. the induced graph of $U^{\prime}$ in $G$ i.e. the subgraph with edge set $E(G[U^{\prime}])=U^{\prime}$ and vertex set $V(G[U^{\prime}])=V(G)$). Given a proper (partial) $k$-edge-colouring $c$ and a set of colours $S\subset[k]$, we let $G[S]$ be the graph containing all edges coloured with a colour from $S$, and all of the vertices incident to an edge coloured with a colour from $S$. Given a proper (partial) $t$-edge-colouring of a graph $G$, we define the set $A(v)$ of available colours at a vertex $v$ to consist of precisely the colours of $[t]$ that no edge incident to $v$ has received.
For a directed graph $G$, any edge denoted by $(uv)\in E(G)$ is oriented from $u$ towards $v$. We denote by $d_{\text{in}}(v)$ (resp. $d_{\text{out}}(v)$) the in-degree of $v$ (resp. out-degree of $v$). Similarly, if $E^{\prime}\subset E(G)$ is a subset of edges and any subgraph $H$, we let $N^{j}_{E^{\prime},\text{in}}(H)$ denote the $j$-hop in-neighbourhood of $H$ in the graph induced by $E^{\prime}$. Similarly, we let $E_{E^{\prime},\text{in}}(H)$ be the edge set containing all in-edges of $H$ also belonging to $E^{\prime}$.
3.1 Chains and shifts
In order to extend a partial edge-colouring, we will have to assign some coloured edge new colours. In order to describe such a process formally, we briefly recall the chain and shift terminology as it was used by for example Bernshteyn [Ber22] (see [Ber22] for a more in-depth treatment).
Let $e_{1},e_{2}\in E(G)$ be two adjacent edges in a graph $G$, and let $c$ be a proper partial colouring of $G$. Then we define a colouring $\operatorname{Shift}(c,e_{1},e_{2})$ by setting:
$\operatorname{Shift}(c,e_{1},e_{2})(e)=\begin{cases}c(e_{2})&\quad\text{if }e=e_{1}\\ \neg&\quad\text{if }e=e_{2}\\ c(e)&\quad\text{if }e\notin\{e_{1},e_{2}\}\\ \end{cases}$
We say that such a pair of adjacent edges $e_{1},e_{2}$ are $c$-shiftable if $c(e_{1})=\neg$ and $c(e_{2})\neq\neg$ and the colouring $\operatorname{Shift}(c,e_{1},e_{2})$ defined above is a proper partial colouring. If the colouring $c$ is clear from the context, we will sometimes leave out this argument. We will only consider shiftable chains.
A chain $C$ of size $k$ is then a set of edges $C=(e_{1},\dots,e_{k})$ such that $e_{i}$ and $e_{i+1}$ are adjacent for all $i$ and $c(e_{i})=\neg$ if and only if $i=1$. For $10\leq j\leq k-$ 1, we can $j$-shift such a chain by performing $\operatorname{Shift}_{j}(c,C)$ defined as:
$\operatorname{Shift}_{0}(c,C)$ $=c$
$\operatorname{Shift}_{i}(c,C)$ $=\operatorname{Shift}(\operatorname{Shift}_{i-1}(c,C),e_{i},e_{i+1})$
We say that $C$ is $c$-shiftable if every pair of edges $e_{i},e_{i+1}$ is $\operatorname{Shift}_{i-1}(c,C)$-shiftable. It is straightforward to check that $j$-shifting such a chain $C$ yields a proper colouring, where the unique uncoloured edge in $C$ is the edge $e_{j+1}$. We say that the chain ends at the edge $e_{k}$ and at the vertex $v$ that is shared by $e_{k}$ and $e_{k-1}$. Note that in the context of simple graphs, this vertex is well defined. We will refer to the act of $k$-shifting a chain of size $k$ as simply shifting the chain. We let $c$ be the pre-shift colouring of $C$ and $\operatorname{Shift}(c,C)$ the post-shift colouring of $C$. We briefly note that from an algorithmic point of view, we can view $\operatorname{Shift}(c,e_{1},e_{2})(e)$ as first uncolouring $e_{2}$, and then assigning $e_{1}$ the colour $c(e_{2})$. This ensures that during all steps of the recolouring, the current colouring of the graph is proper.
Often we wish to find a chain $C$ such that shifting $C$ yields an uncoloured edge $uv$ such that $A(u)\cap A(v)\neq\emptyset$ so that we may colour $uv$ to extend our partial colouring. If this is the case, we say $C$ is an augmenting chain, otherwise we refer to it is a truncated chain. We say that $\kappa$ is available at $uv$ if $\kappa\in A(u)\cap A(v)$.
The initial segment of length $s$ of a chain $C$ is then the chain $C|s=(e_{1},\dots,e_{s})$. One can concatenate two chains $C_{1}=(e_{1},\dots,e_{s_{1}})$ and $C_{2}=(e_{s_{1}},f_{2},\dots,f_{s_{2}})$ to get the chain $C_{1}+C_{2}=(e_{1},\dots,e_{s_{1}},f_{2},\dots,f_{s_{2}})$. Then in order to shift such a chain, we use $\operatorname{Shift}(c,C_{1}+C_{2})=\operatorname{Shift}(\operatorname{Shift}(c,C_{1}),C_{2})$.
A fan chain with center $u$ is a chain $F$ of the form $F=(uw_{1},\dots,uw_{k})$. Here $u$ is the center of the fan, denoted by $\text{center}(F)$. Bernshteyn [Ber22] showed that for such fans to be shiftable chains, we require $c(uw_{i+1})\in A(w_{i})$. We will let the colour of $c(uw_{i+1})$ be the representative available colour at $w_{i}$. We will also pick a representative available colour at $w_{k}$ in $A(w_{k})$. If this representative available colour is also either available at $u$ or if it is the representative available colour for some $w_{j}$ with $j<k$, then we say that the fan is maximal.
A path chain is a shiftable chain of the form $P=(e_{1},\dots,e_{k})$, such that the edges in the chain form a path in the graph. The length of a path chain $P=(e_{1},\dots,e_{k})$ is $k$. A bichromatic path chain is then a $c$-shiftable chain $P=(e_{1},\dots,e_{k})$ that forms a path in $G$ such that $c(e_{1})=\neg$ and the colour of $c(e_{2i})=\kappa_{1}$ for all $1\leq i\leq\lfloor\tfrac{k}{2}\rfloor$ and $c(e_{2i+1})=\kappa_{2}$ for all $1\leq i\leq\lfloor\tfrac{k-1}{2}\rfloor$, and furthermore such that $e_{2}$ is either the first or the last edge in some maximal $(\kappa_{1},\kappa_{2})$-bichromatic path. In order to specify the colours present in the chain, we will sometimes refer to such a chain as a $(\kappa_{1},\kappa_{2})$-bichromatic path chain.
Vizing originally showed how to choose both a fan chain $F$ and a bichromatic path chain $P$ that together form a bigger chain $F+P$ such that $\operatorname{Shift}(c,F+P)$ leaves some edge $uv$ with an available colour. This type of chain has been referred to as a Vizing chain in the literature [Ber22, BD23, Chr23, GP20].
Definition 9.
[…]
We say that $F+P$ ends at the edge $p_{t-1}p_{t}$ and at the vertex $p_{t-1}$. We let $\operatorname{Shift}(c,F+P)$ be the post-shift colouring. We let $|E(P)|$ be the length of $F+P$.
Definition 10.
Given a proper partial colouring $c$ of $G$, an $i$-step Vizing chain is a $c$-shiftable chain of the form $F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i}$ where $F_{j}+P_{j}$ is a Vizing chain for all $j$.
A (multi-step) Vizing chain is augmenting if shifting the chain leaves the final edge of the chain with a colour that is available at both endpoints of the edge. If the chain is not augmenting, it is truncated. The length of a multi-step Vizing chain is the sum of the lengths of the chains $F_{j}+P_{j}$. We will only be considering Vizing chains that are non-overlapping:
Definition 11.
An $i$-step Vizing chain is non-overlapping if every pair of Vizing chains $F_{j}+P_{j}$ on edge $e_{j}$ and $F_{k}+P_{k}$ on edge $e_{k}$ share an edge exactly when $j=k-1$ and the shared edge is $e_{k}$, and if, furthermore, for all $k\leq i$ no edge in the chain $F_{k}+P_{k}=(e_{1},\dots,e_{s})$ is repeated i.e. $e_{j}=e_{h}$ iff $j=h$.
3.2 Algorithms for constructing and extending Vizing chains
In this subsection, we briefly recall some well-known algorithms for constructing Vizing chains. In some cases, one is allowed to make arbitrary choices in the constructions. In order to perform the constructions in a consistent manner, we will order the colours $[s]$ according to the strict ordering $<$ on $\mathbb{N}$. Now, whenever an arbitrary choice between colours (possibly satisfying some conditions) can be made, we will always choose the smallest colour according to $<$. We say that a fan is consistent if the representative available colour $\alpha_{i}$ of each edge $uw_{i}$ is the smallest colour in $A(w_{i})$ subject to the requirements that 1)1)1 ) if $A(u)\cap A(w_{i})\neq\emptyset$, then $\alpha_{i}\in A(u)\cap A(w_{i})$ and 2)2)2 ) else if $j$ is the smallest integer $j<i$ with $\alpha_{j}\in A(w_{i})$ then $\alpha_{i}=\alpha_{j}$. Given a consistent fan-chain $F=(uw_{1},\dots,uw_{k})$, a $(\kappa_{1},\kappa_{2})$-bichromatic path chain $P$ is consistent with $F$ if: $\kappa_{1}\in A(u)$, $\kappa_{2}\in A(w_{k})$ are also chosen as small as possible according to $<$.
The following algorithm will be used to construct the first fan of every Vizing chain. It is a standard result that we have slightly adjusted to have a consistent output.
Lemma 12 ([Ber22, BD23, Viz64]).
Let $G$ be a graph with maximum degree $\Delta$ satisfying $\Delta\leq\Delta_{\max}$. Let $c$ be a proper partial $(\Delta_{\max}+1)$-edge-colouring, and let $e=uv\in E(G)$ be an uncoloured edge. Then there exists a unique consistent maximal fan chain $F=(uw_{1},\dots,uw_{k})$ with center $u$.
Furthermore, $F$ may be computed in $\tilde{O}(\Delta_{\max}^{3})$ time.
Proof.
We first describe how to construct the fan: we set $w_{1}=v$. We pick an available representative colour $\alpha_{1}$ in $A(v)$ for $v$ as follows: if $A(v)\cap A(u)\neq\emptyset$, we let $\alpha_{1}$ be the smallest element of $A(v)\cap A(u)$. In this case, the construction terminates, and we conclude that $F$ is unique. Otherwise, we let $\alpha_{1}$ be the smallest element of $A(v)$. We let $w_{2}$ be the unique neighbour of $u$ with $c(uw_{2})=\alpha_{1}$. In this case, we note that any consistent maximal fan chain must have $(uw_{1},uw_{2})$ as an initial segment.
Next assume that we have inductively constructed a fan chain $(uw_{1},uw_{2},\dots,uw_{j})$ with available representative colours $\alpha_{1},\dots,\alpha_{j-1}$, and that any consistent maximal fan chain on $e$ must have $(uw_{1},uw_{2},\dots,uw_{j})$ as an initial segment. We extend the chain as follows: we pick an available representative colour $\alpha_{j}$ in $A(w_{j})$ for $w_{j}$ as follows:
  1. 1.
    if $A(w_{j})\cap A(u)\neq\emptyset$, we let $\alpha_{j}$ be the smallest element of $A(w_{j})\cap A(u)$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
  2. 2.
    else if there exists a $h<i$ such that $\alpha_{h}\in A(w_{j})$, we let $h^{\prime}$ be the smallest such $h$ and set $\alpha_{j}=\alpha_{h}$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
  3. 3.
    Otherwise, we let $\alpha_{j}$ be the smallest element of $A(w_{j})$. We let $w_{j+1}$ be the unique neighbour of $u$ with $c(uw_{j+1})=\alpha_{j}$. In this case, we note that any consistent maximal fan chain must have $(uw_{1},uw_{2},\dots,uw_{j+1})$ as an initial segment.
The above construction must terminate. Indeed, when $j=d(u)-1$ the third case cannot occur, as $u$ has at most $d(u)-1$ coloured edges incident to it.
As for the furthermore part, we can check 1)1)1 ) in $\tilde{O}(\Delta_{\max}^{2})$ by just checking for each colour in $A(w_{j})$ if it belongs to $A(u)$, we can check 2)2)2 ) in $O(\Delta_{\max}^{2})$ time, since we can check if $\alpha_{k}\in A(w_{j})$ in $O(\Delta_{\max})$ time. Finally, we can do 3)3)3 ) in $\tilde{O}(\Delta_{\max})$ time by simply computing $A(w_{j})$ and extracting the minimum. ∎
Given a consistent maximal fan chain $F=(uw_{1},\dots,uw_{k})$ and a parameter $t\in\mathbb{Z}_{\geq 0}$, we can extend $F$ to a Vizing chain $(F^{\prime}+P)_{t}$, or simply $F^{\prime}+P$ if $t$ is clear from the context, in a consistent manner (note $F^{\prime}$ might not be equal to $F$), as follows: if $\alpha_{k}\in A(u)$, we set $P=(uw_{k})$ and $F^{\prime}=F$. Otherwise, we let $\kappa_{1}$ be the smallest colour in $A(u)$ and $\kappa_{2}=\alpha_{k}$. Let $P^{\prime}=p_{1},p_{2},\dots,p_{t^{\prime}}$ be the unique maximal $(\kappa_{1},\kappa_{2})$-bichromatic path with $p_{1}=w_{k}$. If $t<t^{\prime}-1$, we let $F^{\prime}=F$, and we let $P=(uw_{k},w_{k}p_{2},\dots p_{t}p_{t+1})$. Otherwise, if $t\geq t^{\prime}-1$, there are $3$ cases:
Case 1:
if $p_{t^{\prime}}=w_{i}$ with $i<k$, we let $F^{\prime}=(uw_{1},\dots,uw_{i})$ and we let $P=(uw_{i},w_{i}p_{t^{\prime}-1},p_{t^{\prime}-1}p_{t^{\prime}-2},\dots,p_{2}p_{1})$. Note that in this case $F^{\prime}$ is not a maximal fan.
Case 2:
if $p_{t^{\prime}-1}=w_{i+1}$ and $p_{t^{\prime}}=u$ with $i+1<k$, we let $F^{\prime}=(uw_{1},\dots,uw_{i+1})$ and we let $P=(uw_{i+1},w_{i+1}p_{t^{\prime}-2},p_{t^{\prime}-2}p_{t^{\prime}-3},\dots,p_{2}p_{1})$. Note that in this case $F^{\prime}$ is not a maximal fan.
Case 3:
if $p_{t^{\prime}}\notin\{w_{1},\dots,w_{k}\}$, we let $F^{\prime}=F$ and we let $P=(uw_{k},w_{k}p_{2},p_{2}p_{3},\dots,p_{t^{\prime}-1}p_{t^{\prime}})$.
It is straight-forward to check that we can construct $F^{\prime}+P$ in $O(\Delta_{\max}^{3}+\Delta_{\max}\cdot{}t)$ time, and that the construction is unique. If $e=uv$ is any uncoloured edge and $F$ is the unique maximal fan chain guaranteed by Lemma 12 centered on $u$, then for any given $t$, we say that $(F^{\prime}+P)_{t}$ is a consistent Vizing chain with center $u$. Note that $(F^{\prime}+P)_{t}$ is in fact unique.
Next we will see a different algorithm needed to extend any $i$-step Vizing chain to an $(i+1)$-step Vizing chain in a consistent way. Again we will first construct a fan chain, and then explain how to compute the accompanying bichromatic path chain. We will use the so-called ’second fan lemma’ [Ber22, GP20].
Lemma 13 ([Ber22, GP20]).
Let $G$ be a graph with maximum degree $\Delta$ satisfying $\Delta\leq\Delta_{\max}$. Let $c$ be a proper partial $(\Delta_{\max}+1)$-edge-colouring, and let $e=uv\in E(G)$ be an uncoloured edge. Let $\kappa_{1}\in[\Delta+1]$ be an available colour at $u$ and let $\kappa_{2}\in[\Delta+1]$ be an available colour at $v$. Then there exists a fan chain $uw_{1},\dots,uw_{k}$ with $w_{1}=v$ such that either
  1. 1.
    the representative available colour at $w_{k}$ is also available at $u$.
  2. 2.
    $\kappa_{2}\notin A(u)$ and the fan is maximal subject to the constraint that the colours $\kappa_{1}$ or $\kappa_{2}$ are considered unavailable to every vertex.
  3. 3.
    $w_{k}\neq v$, the representative available colour at $w_{k}$ is $\kappa_{2}$ and no edge $uw_{i}$ is coloured $\kappa_{2}$.
If the representative available colours are chosen as small as possible subject to first 1)1)1 ), then 2)2)2 ), and finally 3)3)3 ) above, then $F$ is unique.
Furthermore, we may compute $F$ in $\tilde{O}(\Delta_{\max}^{3})$ time.
Proof.
We construct a maximal fan almost exactly as in the proof of Lemma 12. Again we begin by setting $w_{1}=v$. We pick an available representative colour $\alpha_{1}$ in $A(v)$ for $v$ as follows: if $A(v)\cap A(u)\neq\emptyset$, we let $\alpha_{1}$ be the smallest element of $A(v)\cap A(u)$. In this case, the construction terminates, and we conclude that $F$ is unique. Otherwise, we let $\alpha_{1}$ be the smallest element of $A(v)\setminus\{\kappa_{2}\}$. Note that since $uv$ is uncoloured, $|A(v)|\geq 2$. We let $w_{2}$ be the unique neighbour of $u$ with $c(uw_{2})=\alpha_{1}$. In this case, we note that any consistent maximal fan chain must have $(uw_{1},uw_{2})$ as an initial segment.
Next assume that we have inductively constructed a fan chain $(uw_{1},uw_{2},\dots,uw_{j})$ with available representative colours $\alpha_{1},\dots,\alpha_{j-1}$, and that any consistent maximal fan chain on $e$ must have $(uw_{1},uw_{2},\dots,uw_{j})$ as an initial segment. We extend the chain as follows: we pick an available representative colour $\alpha_{j}$ in $A(w_{j})$ for $w_{j}$ as follows:
  1. 1.
    if $A(w_{j})\cap A(u)\neq\emptyset$, we let $\alpha_{j}$ be the smallest element of $A(w_{j})\cap A(u)$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
  2. 2.
    else if there exists a $h<i$ such that $\alpha_{h}\in A(w_{j})$, we let $h^{\prime}$ be the smallest such $h$ and set $\alpha_{j}=\alpha_{h}$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
  3. 3.
    else if $|A(w_{j})\setminus\{\kappa_{2}\}|\geq 1$, we let $\alpha_{j}$ be the smallest element of $A(w_{j})\setminus\{\kappa_{2}\}$. We let $w_{j+1}$ be the unique neighbour of $u$ with $c(uw_{j+1})=\alpha_{j}$. In this case, we note that any consistent maximal fan chain must have $(uw_{1},uw_{2},\dots,uw_{j+1})$ as an initial segment.
  4. 4.
    Otherwise, we set $\alpha_{j}=\kappa_{2}$ and conclude the construction.
The above construction must terminate. Indeed, when $j=d(u)-1$ the third case cannot occur, as $u$ has at most $d(u)-1$ coloured edges incident to it.
Similarly to Lemma 12, we can construct the fan in $\tilde{O}(\Delta_{\max}^{3})$ time. ∎
Now given some $i$-step Vizing chain $F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i}$ and some parameter $t\in\mathbb{Z}_{\geq 0}$, we can extend it in a consistent manner by applying Lemma 13 to the colouring $\operatorname{Shift}(c,F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i})$ and the final edge $p_{t_{i}-1}p_{t_{i}}=u^{\prime}v^{\prime}$ to get a unique fan $F=(u^{\prime}w_{1}^{\prime},u^{\prime}w_{2}^{\prime},\dots,u^{\prime}w_{k}^{\prime})$. Let $P_{i}$ be $(\kappa_{1},\kappa_{2})$-bichromatic with $c(u^{\prime}v^{\prime})=\kappa_{2}$, then we apply Lemma 13 with $\kappa_{1}$ and $\kappa_{2}$. If we are in Case $1$, we set $F_{i+1}=F$ and $P_{i+1}=u^{\prime}w_{k}^{\prime}$. If we are in Case $2$, we set $F_{i+1}=F$ and let $\tau_{1}\in A(u^{\prime})\setminus\{\kappa_{1},\kappa_{2}\}$ and $\tau_{2}=\alpha_{k}$ be the smallest available colours. Note that $|A(u^{\prime})|\geq 2$ in the shifted colouring, and that $\kappa_{2}\notin A(u)$ by assumption. Then we let $P^{\prime}=p^{\prime}_{1}p^{\prime}_{2}\dots p^{\prime}_{t^{\prime}}$ be the maximal $(\tau_{1},\tau_{2})$-bichromatic path beginning at $p^{\prime}_{1}=w_{k}^{\prime}$. If $t<t^{\prime}-1$, we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t}p_{t+1})$, otherwise we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t^{\prime}-1}p_{t^{\prime}})$. Otherwise, we are in Case 3)3)3 ). Again we set $F_{i+1}=F$, but this time we let $P^{\prime}=p^{\prime}_{1}p^{\prime}_{2}\dots p^{\prime}_{t^{\prime}}$ be the maximal $(\kappa_{1},\kappa_{2})$-bichromatic path beginning at $p^{\prime}_{1}=w_{k}^{\prime}$. If $t<t^{\prime}-1$, we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w^{\prime}_{k}p_{2},\dots p_{t}p_{t+1})$, otherwise we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t^{\prime}-1}p_{t^{\prime}})$.
Similarly to before, we can perform the above extension in $\tilde{O}(\Delta_{\max}^{3}+\Delta_{\max}\cdot{}t)$ time. We will say that we extend the Vizing chain via Lemma 13 in a consistent manner. Note that if every extension in a Vizing chain $F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i}$ is consistent, then $P_{j+1}$ will never overlap $P_{j}$ in any edges for all $j$. Indeed, the colours of $P_{j+1}$ are either disjoint from those of $P_{j}$ or identical to those of $P_{j}$. In the latter case, $P_{j+1}$ and $P_{j}$ belong to different components in the shifted colouring.
In this paper, we will only be working with Vizing chains satisfying certain restrictions. The restrictions we will impose are as follows:
Definition 14.
Let $VC=F_{1}+P_{1}+\dots+F_{j}+P_{j}$ be a non-overlapping multi-step Vizing chain. We say that $VC$ is consistent if the following conditions hold:
  • $F_{1}+P_{1}$ is constructed in a consistent manner (Lemma 12 and following discussion).
  • For $i\geq 2$ we have that $F_{i}+P_{i}$ is constructed in a consistent manner (Lemma 13 and following discussion).
[…]
  • For all $j$, we have that the maximal bichromatic path containing $P_{j}$ under $c$ is different from the maximal bichromatic path containing $P_{j+1}$ under $c$.
3.3 Path-stepping processes
We will now consider a looser process than the one used to construct Vizing chains. Beginning from some vertex $w$, we let one step be the process of first moving to a neighbour of $w$, say $y$, and then following some bichromatic path $P$ having $y$ as an endpoint. As with the multi-step Vizing chains we are allowed to truncate the bichromatic path at any point to end up at some new vertex $w^{\prime}$. We may then iteratively repeat this process at $w^{\prime}$. If we do so $i$ times, we say that we run an $i$-step process starting from $w_{1}$. We denote by $w_{1}+y_{1}+\hat{P}_{1}+w_{2}+y_{2}+\hat{P}_{2}+\dots w_{i}+y_{i}+\hat{P}_{i}$ the $i$-step process beginning at $w$, that first jumps to the neighbour $y_{1}$ of $w$, then follows a truncated maximal bichromatic path $P_{1}$ to end up at $w_{2}$, and then jumps to the neighbour $y_{2}$ of $w_{2}$ and follows a truncated maximal bichromatic path $P_{2}$ and so on and so forth. We will say that the process is non-overlapping if for all $k,j$ we have that $V(P_{k})\cap V(P_{j})\neq\emptyset$ implies that $|k-j|\leq 1$, if $E(P_{k})\cap E(P_{j})\neq\emptyset$ implies $k=j$, if $V(P_{k})\cap N^{2}(w_{j})\neq\emptyset$ for $k\geq j$ implies that $k=j$, and, finally, if the maximal bichromatic path $Q_{j}$ containing $P_{j}$ is equal to the maximal bichromatic path $Q_{k}$ containing $P_{k}$, then $j=k$. In other words, we will only allow a bichromatic path to intersect the 2-hop-neighbourhood of the immediately preceding path, and we will not allow it to intersect the edge set of any previous paths. If $|P_{j}|+1\leq\ell$ for all $j$, we say that the process has step length at most $\ell$. If $P_{i}$ is a maximal bichromatic path, we say that the $i$-step process is augmenting.
Maintaining information about this type of process is useful, as it allows us to accurately represent where consistent Vizing chains might go, while at the same time avoiding the need to worry about exactly which extensions are possible at what times. In particular, we observe that any consistent and non-overlapping $i$-step Vizing chain $F_{1}+P_{1}+\dots+F_{i}+P_{i}$ where $F_{1}$ is centered at $w$ naturally induces an $i$-step process on $w$, by letting $w_{j}=\text{center}(F_{j})$, letting $y_{j}$ be the second point of $P_{j}$, and letting $\hat{P}_{j}$ be the the bichromatic sub-path of the path $P_{j}$ excluding only the first edge of the chain $P_{j}$. To make this observation more precise, we will define and consider semi-consistent processes. Here, we will put some further restrictions on the processes so that they model consistent Vizing chains exactly.
Given an $i$-step process $w_{1}+y_{1}+P_{1}+w_{2}+y_{2}+P_{2}+\dots w_{i}+y_{i}+P_{i}$, we say that the extension $w_{j}+y_{j}+P_{j}$ for $j\geq 2$ is consistent with Lemma 13, if the following is the case: consider the colouring $c^{\prime}$ obtained from $c$ as follows: suppose $P_{j-1}=p_{1}p_{2}\dots p_{t}$ is $(\kappa_{1},\kappa_{2})$-bichromatic with $c(p_{t-1}p_{t})=\kappa_{1}$. Suppose the $P=p_{1}p_{2}\dots p_{t}p_{t+1}\dots p_{s}$ is a maximal bichromatic path containing $P_{j-1}$. Note that if $s=t$, then $P_{j-1}$ is augmenting, and it cannot be extended consistently.
We obtain $c^{\prime}$ by first un-colouring the edge $p_{t}p_{t+1}$ i.e. $c^{\prime}(p_{t}p_{t+1})=\neg$, and then setting $c^{\prime}(p_{k}p_{k+1})=\kappa_{2}$ if $c(p_{k}p_{k+1})=\kappa_{1}$ and vice versa for $k<t-1$ setting $c^{\prime}(p_{k}p_{k+1})=\kappa_{1}$ if $c(p_{k}p_{k+1})=\kappa_{2}$. It is straight-forward to check that if $c$ is a proper $(\Delta_{\max}+1)$-edge-colouring, then so is $c^{\prime}$. Now, apply Lemma 13 to $p_{t-1}p_{t}$ with the colouring $c^{\prime}$, $u=p_{t-1}$, $v=p_{t}$, and the colours $\kappa_{1}$ and $\kappa_{2}$ as before. By fixing a length $t^{\prime}=|P_{i}|$, the lemma together with the discussion that follows it produces a fan chain $F$ and a bichromatic-path chain $Q=(q_{0}q_{1},\dots,q_{\hat{t}}q_{\hat{t}+1})$ with $\hat{t}\leq t^{\prime}$ such that $F+Q$ is a $1$-step Vizing chain. We say the extension $w_{j}+y_{j}+P_{j}$ for $j\geq 2$ is consistent with Lemma 13 if $w_{j}=\text{center}(F_{j})=q_{0}$, and if $y_{j}=q_{1}$, and finally $P_{j}=q_{1},q_{2},\dots,q_{t^{\prime}}$. If $\hat{t}=t^{\prime}-1$, then $P_{j}$ is augmenting. If $\hat{t}<t^{\prime}-1$, then the extension is not consistent.
Next, we define a semi-consistent process:
Definition 15.
Let $C=w_{1}+y_{1}+P_{1}+w_{2}+y_{2}+P_{2}+\dots w_{i}+y_{i}+P_{i}$ be a non-overlapping $i$-step process beginning at $w_{1}$. We say that $C$ is semi-consistent if the following holds:
  • For $j\geq 2$ we have that $w_{j},y_{j}$ and $P_{j}$ is an extension consistent with Lemma 13.
  • We have that $w_{j}$ is disjoint from the $2$-hop neighbourhood of $w_{1}+y_{1}+P_{1}+\dots w_{j-2}+y_{j-2}+P_{j-2}+w_{j-1}$ for all $j$.
  • For all $j$, we have that $P_{j}$ is disjoint from the $2$-hop neighbourhood of $w_{k}$ for all $k\leq j-1$ and from $P_{k}$ for all $k\leq j-2$.
We observe that for any $j<i$ the length of $P_{j}$ is at least $3$. Indeed, otherwise Item 2 in Definition 15 fails. The following Lemma is immediate:
Lemma 16.
Let $G$ be a graph endowed with a proper (partial) edge-colouring $c$. Let $\mathcal{SCP}$ be the set of semi-consistent processes on $(G,c)$, and let $\mathcal{NOP}$ be the set of non-overlapping processes on $(G,c)$. Then we have:
$\mathcal{SCP}\subset\mathcal{NOP}$
Semi-consistent processes are in correspondence with consistent Vizing chains in the following sense:
Lemma 17.
There is a mapping $\gamma$ from the set of consistent $i$-step Vizing chains to the set of semi-consistent $i$-step processes. In particular, if a consistent $i$-step Vizing chains augmenting, so is its image, and similarly if a semi-consistent $i$-step process is augmenting, so is every Vizing chain in its pre-image.
Proof.
Let $VC=F_{1}+Q_{1}+\dots+F_{i}+Q_{i}$ be any consistent $i$-step Vizing chain. Then we will let $VC=F_{1}+Q_{1}+\dots+F_{i}+Q_{i}$ correspond to the $i$-step process $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$, where $w_{j}=\text{center}(F_{j})$, if $Q_{j}=(q_{0}q_{1},\dots,q_{t_{j}+1})$, then $y_{j}=q_{1}$ and $P_{j}=q_{1},q_{2}\dots,q_{t_{j}}$.
Next we show that $C$ is semi-consistent. First of all, we need to check that $C$ is non-overlapping. This follows from the fact that $VC$ is consistent, which puts precisely the same constraints on the $P_{j}$’s due Definition 14 as is required for $C$ to be non-overlapping. The first item of Definition 15 is satisfied since the second item of Definition 14 holds. This is the case since no edge outside of $P_{j-1}$ in the two-hop neighbourhood of the center of $F_{j}$ has changed colour before Lemma 13 is applied. This means that applying it directly to $P_{j-1}$ yields the same result as using it to extend the Vizing chain. Finally, items 3 and 4 of Definition 14 are equivalent to the last two items of Definition 15.
Furthermore, observe that solely the length of the maximal path containing $P_{i}$ determines whether $C$ and $VC$ are augmenting and so $C$ is augmenting if and only if $VC$ is.
The other direction is symmetrical. ∎
If a semi-consistent process is the image of a consistent $i$-step Vizing chain, we say that the process is consistent. Finally, we note that under the right conditions, we can use semi-consistent processes to build consistent Vizing chains:
Lemma 18.
Let $G$ be a graph with maximum degree $\Delta\leq\Delta_{\max}$, and let $c$ be a proper $(\Delta_{\max}+1)$-edge-colouring of $G$. Let $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ be an augmenting semi-consistent $i$-step process with step-length $\leq\ell$. Let $e=w_{1}v$ be an un-coloured edge. Suppose that applying Lemma 12 and the following discussion yields a consistent $1$-step Vizing chain $(F+Q_{1})_{\ell}$ such that
[…]
  • $y_{1}=x_{k}$ with $F=(w_{1}x_{1},w_{1}x_{2},\dots w_{1}x_{k}$
  • $P_{1}=q_{1}q_{2}\dots q_{\ell}$ with $Q=(q_{0}q_{1},q_{1}q_{2},\dots,q_{\ell}q_{\ell+1})$
Then there is an augmenting consistent $i$-step Vizing chain on $e$ with the same step length as $C$.
Proof.
We will construct the augmenting consistent Vizing chain on $e$, say $VC=F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{i}+Q_{i}$ as follows: we set $F_{1}=F$ and $Q_{1}=Q$. Then we extend $F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j}$ by applying Lemma 13 and the following discussion making sure that the length of $F_{j+1}+Q_{j}$ is equal to $t_{j}+2$ with $t_{j}=|P_{j}|$.
We now observe that $\gamma(VC)=C$. Indeed, by construction every extension results in $Q_{j}=(q_{0,j}q_{1,j},\dots,q_{t_{j}+1,j}q_{t_{j}+2,j})$ such that $P_{j}=q_{1,j}q_{2,j}\dots q_{t_{j}+1,j}$. This follows from the fact that since $C$ is semi-consistent, the extension performed on $P_{j}$ uses a colouring $c^{\prime}$ which only differs from the colouring $\operatorname{Shift}(c,F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j})$ outside of the $2$-hop neighbourhood of $w_{j}$. Indeed, $c^{\prime}$ and $\operatorname{Shift}(c,F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j})$ agree on the colours along $P_{j}$ by construction, and $w_{j}+1$ is disjoint from the two-hop neighbourhood of $C=w_{1}+y_{1}+P_{1}+\dots+w_{j-1}+y_{j-1}+P_{j-1}+w_{j}$ also by construction. Since $C$ is semi-consistent, the extension ensures that $Q_{j}=(q_{0,j}q_{1,j},\dots,q_{t_{j}+1,j}q_{t_{j}+2,j})$ such that $P_{j}=q_{1,j}q_{2,j}\dots q_{t_{j}+1,j}$.
It now follows that $\gamma(VC)=C$, and therefore Lemma 17 now guarantees that $VC$ is also augmenting. ∎
4 Dynamic t-splitters
Given a graph $G$ equipped with any (not necessarily proper) edge-colouring $c$, we let $d_{i}(u)$ be the degree of $u$ in the graph induced in $G$ by all edges coloured with the colour $i$. We let $\tilde{d}=\max_{\kappa}\max_{v}d_{\kappa}(v)$. Given two disjoint subsets of vertices $A,B\subset V(G)$, we let $G[A,B]$, be the bipartite subgraph of $G$ on vertex set $A\cup B$ containing exactly the edges going from a vertex in $A$ to a vertex in $B$. We let $s_{i}(v)=\max\{d_{i}(v)-\frac{\Delta_{\max}}{t},0\}$ be the surplus of $i$-coloured edges around $v$. Finally, we let $L:E(G)\rightarrow\mathbb{Z}$ be a function that maps $uv\mapsto s_{c(uv)}(u)+s_{c(uv)}(v)$, and we set $\max L(e)=\tilde{L}$.
In this section, we will show Theorem 3, which we restate below for convenience:
Theorem 19 (Identical to Theorem 3).
Let $G$ be a dynamic graph, and suppose $\varepsilon>0$, $t\in\mathbb{N}$, and an upper bound $\Delta_{\max}$ on the maximum degree of $G$ through the entire update sequence is given. Assume that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then there is a dynamic algorithm that computes a $t$-splitter of $G$ with an amortised update-time of $O(t^{3}\varepsilon^{-4}\log^{3}m)$.
In this entire section, we will work with two parameters $\Delta_{\max}$, which we assume is a known upper-bound on the maximum degree of some dynamic graph $G$, as well as a parameter $\eta\geq 4$. Note that it is important that $\eta\geq 1$, but we will only apply the theory below, when this is the case. We will analyse a colouring $c$ satisfying the following invariant:
Invariant 20.
For all $uv\in E(G)$ and all $i\in[t]$, we have that
$L(uv)=s_{c(uv)}(u)+s_{c(uv)}(v)\leq\eta+s_{i}(u)+s_{i}(v)$
We first show that if $1\leq\eta\leq\tfrac{\varepsilon^{2}}{128\log n}\tfrac{\Delta_{\max}}{t}$, then $\tilde{d}\leq(1+\varepsilon)\frac{\Delta_{\max}}{t}$. We note already now that we will only apply the above Theorem, when $\tfrac{\Delta_{\max}}{t}=\Omega(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$, so the assumption that $\eta\geq 1$ is not dangerous.
4.1 Analysing the Local Condition
We first show the following structural theorem:
Theorem 21.
Let $0<\varepsilon\leq\frac{1}{2}$ and let $G$ be equipped with a $t$-colouring $c$ satisfying Invariant 20. Let $1\leq\eta\leq\tfrac{\varepsilon^{2}}{128\log n}\tfrac{\Delta_{\max}}{t}$. Then it holds that $\tilde{d}\leq(1+\varepsilon)\frac{\Delta_{\max}}{t}$.
Before we give this proof, we first show some helpful lemmas. The first lemma shows that if an edge $e$ has a large surplus, then any endpoint of $e$ has a large surplus in some colour.
Lemma 22.
Suppose Invariant 20 holds. Suppose furthermore that $e=uv$ has $L(uv)=D$. Then there exists a colour $\kappa$ such that $s_{\kappa}(u)\geq D-\eta$.
Proof.
Consider $v$. Since $d(v)\leq\Delta_{\max}$, it follows by the pigeon-hole principle that there exists some colour $\kappa^{\prime}$ for which $d_{\kappa^{\prime}}(v)\leq\frac{\Delta}{t}$. Since we assumed Invariant 20 holds, we have that
$D\leq\eta+s_{\kappa^{\prime}}(u)+s_{\kappa^{\prime}}(v)=\eta+s_{\kappa^{\prime}}(u)$
Hence, we deduce that $D-\eta\leq s_{\kappa^{\prime}}(u)$, and we are done by setting $\kappa=\kappa^{\prime}$. ∎
We let $\tilde{L}:=\max_{e\in E(G)}L(e)$. The next lemma shows how to inductively determine a large set of bipartite subgraphs containing vertices with large surpluses.
Lemma 23.
Suppose Invariant 20 holds, and let $\mu\in\mathbb{Z}$ be a parameter. Then there exists two colours $\kappa_{x}$ and $\kappa_{y}$, as well as two family of sets $(X_{i})_{i=1}^{\mu}$ and $(Y_{i})_{i=1}^{\mu}$ such that for all $i$, we have $X_{i},Y_{i}\subset V(G)$ and that the following two conditions hold:
$\forall u\in X_{i}:s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)$ $\geq\tilde{L}-(1+2(i-1))\eta$ (1)
$\forall v\in Y_{i}:s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)$ $\geq\tilde{L}-2i\eta$ (2)
Proof.
We first determine $\kappa_{x}$ and $\kappa_{y}$. Then we construct the sets inductively. Pick an arbitrary edge $uv$ achieving $L(uv)=\max L(e)=\tilde{L}$. Now we may apply Lemma 22 to see that there exists a colour $\kappa_{x}$ such that $d_{\kappa_{x}}(u)\geq\frac{\Delta_{\max}}{t}+\tilde{L}-\eta$. By the pigeon-hole principle there is some colour $\kappa_{y}$ incident to $u$ such that $d_{\kappa_{y}}(u)\leq\frac{\Delta_{\max}}{t}$.
With $\kappa_{x}$ and $\kappa_{y}$ defined, we construct the family of sets inductively. Initially, we set $X_{1}=\{u\}$ and $Y_{1}=N_{G[\kappa_{x}]}(u)-\{u\}$. We confirm that the required conditions hold: $(1)$ is clear from our choice of $\kappa_{x}$ and $\kappa_{y}$. We get $(2)$ by observing that by Invariant 20, we find that for any $v\in Y_{1}$, we have:
$s_{\kappa_{y}}(v)+s_{\kappa_{y}}(u)+\eta\geq s_{\kappa_{x}}(v)+s_{\kappa_{x}}(u)$
Hence:
$s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)$ $\geq s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)-\eta$
$\geq\tilde{L}-2\eta$
by above.
For the inductive step, we now set $X_{2}=N_{G[\kappa_{y}]}(Y_{1})-Y_{1}$, and $Y_{2}=N_{G[\kappa_{x}]}(X_{2})-X_{2}$. Again we confirm the two conditions:
  • $(1)$: Consider $u\in X_{i}$. We have $\forall v\in Y_{i-1}$ that $s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)\geq\tilde{L}-2(i-1)\eta$ by the induction hypothesis. By applying Invariant 20, we find that
    $s_{\kappa_{x}}(v)+s_{\kappa_{x}}(u)+\eta\geq s_{\kappa_{y}}(v)+s_{\kappa_{y}}(u)$
    since $c(uv)=\kappa_{y}$. Hence:
    $s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)$ $\geq s_{\kappa_{x}}(v)-s_{\kappa_{y}}(v)-\eta$
    $\geq\tilde{L}-(1+2(i-1))\eta$
    as before.
  • $(2)$ is symmetric: consider any $v\in Y_{i}$. We have $\forall u\in X_{i}$ that $s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)\geq\tilde{L}-(1+2(i-1))\eta$ by induction. By applying Invariant 20, we find that
    $s_{\kappa_{y}}(v)+s_{\kappa_{y}}(u)+\eta\geq s_{\kappa_{x}}(v)+s_{\kappa_{x}}(u)$
    since $c(uv)=\kappa_{x}$. Hence:
    $s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)$ $\geq s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)-\eta$
    $\geq\tilde{L}-2i\eta$
    as before. Note that the above argument can be repeated ad infinitum, and so in particular also $\mu$ times.
We are now ready to show Theorem 21:
Proof of Theorem 21.
Suppose for contradiction that $\tilde{d}>(1+\varepsilon)\frac{\Delta}{t}$. We will show that this implies the existence of a family of bipartite subgraphs $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}=\{(P_{j},Q_{j}))\}_{j=1}^{10\log(n)\varepsilon^{-1}}$ of $G$ satisfying that $|Q_{j}|\geq(1+\frac{\varepsilon}{4})^{j}$.
This implies that for some $j\leq 10\log(n)\varepsilon^{-1}$, we have that
$|Q_{j}|\geq(1+\frac{\varepsilon}{4})^{10\log(n)\varepsilon^{-1}}\geq e^{\frac{10}{8}\log n}>n$
where we used the standard in-equality $1-x\leq e^{-x}$ which holds for all real $x$. In particular, it follows that for all $x\leq 1/2$, we have that $e^{x}\leq\frac{1}{1-x}\leq 1+\frac{x}{1-x}\leq 1+2x$. This contradicts the fact that we also must have $\mathbb{B}_{j}\subset G$ and so $|V(\mathbb{B}_{j})|\leq|V(G)|=n$.
In order to construct $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}$, we invoke Lemma 23 with $\mu=\lceil 10\log(n)\varepsilon^{-1}\rceil$ to get sets $(X_{i})_{i=1}^{\mu}$ and $(Y_{i})_{i=1}^{\mu}$ and colours $\kappa_{x}$ and $\kappa_{y}$ satisfying the condition of the lemma. Then for appropriate values of $i$, we set
$\mathbb{B}_{2i-1}=(G[\kappa_{x}])[X_{i},Y_{i}]$
with $P_{2i-1}=X_{i}$ and $Q_{2i-1}=Y_{i}$ as well as
$\mathbb{B}_{2i}=(G[\kappa_{y}])[Y_{i},X_{i+1}]$
with $P_{2i}=Y_{i}$ and $Q_{2i}=X_{i+1}$ These graphs are clearly bipartite by construction, so all we need to show is that $Q_{i}$ must expand. We show this by induction on $i$. The base case is is as follows: since we have tacitly assumed that $\eta\geq 1$ it follows that $\tilde{L}-\eta$ is large, say at least $2$. In particular, we have that $|Q_{1}|=|Y_{1}|\geq\tilde{L}-\eta\geq 2$.
Instead of showing that $Q_{2}$ is also sufficiently large, we proceed to the induction step directly. We will first show that if $Q_{2i-1}\geq(1+\frac{\varepsilon}{4})^{2i-1}$, then $Q_{2i}\geq(1+\frac{\varepsilon}{4})^{2i}$. Then we show that if $Q_{2i}\geq(1+\frac{\varepsilon}{4})^{i}$ , then $Q_{2i+1}\geq(1+\frac{\varepsilon}{4})^{2i+1}$. In this way, we only need to show the base case $|Q_{1}|=|Y_{1}|\geq\tilde{L}-\eta\geq 2$, as we have already done.
So for the inductive step, we first consider $\mathbb{B}_{2i}$. We note that, by construction, we have that $\forall u\in X_{i+1}$
$s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)\geq\tilde{L}-(1+2i)\eta$
similarly for all $v\in Y_{i}$ we have
$s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)\geq\tilde{L}-2i\eta$
This implies that every vertex $v\in Y_{i}$ has $s_{\kappa_{y}}(v)\geq\tilde{L}-2i\eta$ and that every vertex $u\in X_{i+1}$ has $s_{\kappa_{y}}(u)\leq(1+2i)\eta$. Since the graph is bipartite, we have that
$\sum\limits_{u\in X_{i+1}}d_{\kappa_{y}}(u)\geq\sum\limits_{v\in Y_{i}}d_{\kappa_{y}}(v)$
Applying the above, we find that
$((1+2i)\eta+\frac{\Delta}{t})|X_{i+1}|\geq\sum\limits_{u\in X_{i+1}}d_{\kappa_{y}}(u)\geq\sum\limits_{v\in Y_{i}}d_{\kappa_{y}}(v)\geq(\tilde{L}-2i\eta+\frac{\Delta}{t})|Y_{i}|$
Since $\tilde{d}\leq\frac{\Delta}{t}+\tilde{L}$ by definition and since $\eta\frac{22\log(n)}{\varepsilon}\leq\frac{\varepsilon\Delta}{4t}$ by choice of $\eta$, we have that
$\tilde{L}+\frac{\Delta}{t}-\eta\frac{22\log(n)}{\varepsilon}\geq\tilde{d}-\eta\frac{22\log(n)}{\varepsilon}\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}$
and that
$\frac{\Delta}{t}+\eta\frac{22\log(n)}{\varepsilon}\leq(1+\frac{\varepsilon}{4})\frac{\Delta}{t}$
Hence we can write
$(1+\frac{\varepsilon}{4})\frac{\Delta}{t}|X_{i+1}|\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}|Y_{i}|$
Therefore, we find that
$|X_{i+1}|$ $\geq\frac{(1+\frac{3\varepsilon}{4})}{(1+\frac{\varepsilon}{4})}|Y_{i}|$
$\geq(1-\frac{\varepsilon}{4})(1+\frac{3\varepsilon}{4})|Y_{i}|$
$\geq(1+\frac{3\varepsilon}{4}-2\frac{\varepsilon}{4}))|Y_{i}|$
$\geq(1+\frac{\varepsilon}{4})|Y_{i}|$
since $Q_{2i}=X_{i+1}$ and $Q_{2i-1}=Y_{i}$, applying the induction hypothesis yields the first case. The second case is symmetric, but we include it for completeness.
Now consider $\mathbb{B}_{2i+1}$. We note that, by construction, we have that $\forall u\in X_{i+1}$
$s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)\geq\tilde{L}-(1+2i)\eta$
similarly for all $v\in Y_{i+1}$ we have
$s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)\geq\tilde{L}-2(i+1)\eta$
This implies that every vertex $v\in Y_{i}$ has $s_{\kappa_{x}}(v)\leq 2(i+1)\eta$ and that every vertex $u\in X_{i+1}$ has $s_{\kappa_{x}}(u)\geq\tilde{L}-(1+2i)\eta$. Since the graph is bipartite, we have that
$\sum\limits_{v\in Y_{i+1}}d_{\kappa_{x}}(v)\geq\sum\limits_{u\in X_{i+1}}d_{\kappa_{x}}(u)$
Applying the above, we find that
$(2(i+1)\eta+\frac{\Delta}{t})|Y_{i+1}|\geq\sum\limits_{u\in X_{i+1}}d_{\kappa_{y}}(u)\geq\sum\limits_{v\in Y_{i}}d_{\kappa_{y}}(v)\geq(\tilde{L}-(1+2i)\eta+\frac{\Delta}{t})|X_{i+1}|$
Since $\tilde{d}\leq\frac{\Delta}{t}+\tilde{L}$ by definition and since $\eta\frac{22\log(n)}{\varepsilon}\leq\frac{\varepsilon\Delta}{4t}$ by choice of $\eta$, we have similarly to before that
$\tilde{L}+\frac{\Delta}{t}-\eta\frac{22\log(n)}{\varepsilon}\geq\tilde{d}-\eta\frac{22\log(n)}{\varepsilon}\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}$
and that
$\frac{\Delta}{t}+\eta\frac{22\log(n)}{\varepsilon}\leq(1+\frac{\varepsilon}{4})\frac{\Delta}{t}$
So we can write
$(1+\frac{\varepsilon}{4})\frac{\Delta}{t}|Y_{i+1}|\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}|X_{i+1}|$
Therefore, we find, as before, that
$|Y_{i+1}|$ $\geq(1+\frac{\varepsilon}{4})|X_{i+1}|$
since $Q_{2i+1}=Y_{i+1}$ and $Q_{2i}=X_{i+1}$, applying the induction hypothesis thus yields the second case. ∎
Note that Theorem 21 shows that any colouring satisfying the local condition is a $t$-degree splitting colouring.
4.2 Dynamically Maintaining Invariant 20
Again, we assume we have access to some upper bound $\Delta_{\max}$ on the maximum degree over the entire update sequence, and some input parameter $t$. We further require that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. We will set $\eta=\lfloor\tfrac{\varepsilon^{2}}{128\log n}\rfloor\tfrac{\Delta_{\max}}{t}$, and note that $\eta\geq 10$ always due to our assumptions.
We will maintain the following data structures. At each vertex $u$, we will store:
  • The surplus for each colour in an index-able array $SP(u)$.
  • For each colour $\kappa$, we store a BBST, $MSD_{\kappa}(u)$, containing all estimated surplus differences of the form:
    $\tilde{s}_{c(wx)}(w)+\tilde{s}_{c(wx)}(x)-(\tilde{s}_{\kappa}(w)+\tilde{s}_{\kappa}(x))$
  • For each colour $\kappa$, we store a doubly-linked-list $LL(u,\kappa)$ containing all edges coloured $\kappa$ incident to $u$.
  • For each colour $\kappa$, we store a counter $\iota(u,\kappa)\in[d_{k}]$ and a pointer to the $\iota(u,\kappa)^{\text{th}}$ element of $LL(u,\kappa)$.
For each edge $e$, we store a BBST over the estimated surpluses for each colour: $T^{s}_{e}$.
The algorithm now works as follows (see below for pseudo-code). Whenever an edge $e=uv$ is inserted, we calculate the surplus $s_{\kappa}(u)+s_{\kappa}(v)$ for all choices of $\kappa$, and assign $e$ an arbitrary colour $\kappa^{\prime}$ with minimum surplus. We also initialize all of $e$’s data structures, and update its contributions to $u$ and $v$’s data structures. Then, we insert $e$ into the linked-lists $LL(u,\kappa^{\prime})$ and $LL(v,\kappa^{\prime})$, ensuring that $e$ is inserted just before position $\iota(u,\kappa^{\prime})$ in $LL(u,\kappa^{\prime})$ resp.  $\iota(v,\kappa^{\prime})$ in $LL(v,\kappa^{\prime})$. Then we update data structures for other edges $f$ around $w\in\{u,v\}$ in a round-robin fashion: starting with the edge stored at position $\iota(w,\kappa^{\prime})$ in $LL(w,\kappa^{\prime})$, we recompute the contents of the next $\lceil 500\tfrac{\Delta_{\max}}{\eta}\rceil$ edges in $LL(w,\kappa^{\prime})$ and update $\iota(w,\kappa^{\prime})=\iota(w,\kappa^{\prime})+\lceil 500\tfrac{\Delta_{\max}}{\eta}\rceil\mod d_{\kappa^{\prime}}(w)$. Then, we check for all colours if any violations have arisen. If so, we push the endpoint and the colour with too large a surplus difference, unto a queue $Q$. After this, we process $Q$ one pair at the time by recolouring with a similar procedure as above, possibly pushing new violating pairs unto $Q$. Note that to get the edge to recolour, we extract the observed minimum and check if its surplus difference is still large enough. If yes, we process it by recolouring it, if no, we conclude our processing of this element of $Q$. Deletions are symmetric. Here, we have to remember to remove $e$ from any doubly-linked-lists containing it, and to re-adjust the relevant $\iota(\cdot{},\cdot{})$ by decrementing them by $1$, if $e$’s position in $LL(\cdot{},\cdot{})$ was before the element pointed to by $\iota(\cdot{},\cdot{})$.
Note that these running times can probably be optimised with respect to $t$, but since any such optimisations are likely to only change constant factors hidden by big-Oh notation anyways in the final edge-colouring algorithm, we have not attempted them here. Such optimisation, however, would yield a better $t$-splitter.
We update $T_{e}^{s}$ be calculating the surplus for each colour of $e$ and inserting it into a search tree. We then update (deleting the old, and inserting the new) the surplus difference for $e$ into $MSD_{\kappa}(w)$ for all colours $\kappa$ and all endpoints $w$ of $e$. These data structures always contain correct estimates immediately after being updated. Hereafter they might slowly deteriorate in quality, until they are updated again. We will later show that this deterioration is small enough to be handled efficiently. We will denote by $\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v)$ the estimate of a surplus stored in a data structure and by $s_{\kappa}(u)+s_{\kappa}(v)$ the actual value of a surplus. In particular, we will denote by $\Delta\tilde{s}_{\kappa}(w)$ the maximum estimated surplus difference stored in $MSD_{\kappa}(w)$:
$\Delta\tilde{s}_{\kappa}(w)=\max_{x\in N(w)}\tilde{s}_{c(wx)}(w)+\tilde{s}_{c(wx)}(x)-(\tilde{s}_{\kappa}(w)+\tilde{s}_{\kappa}(x))$
Note that this value can be found by querying $MSD_{\kappa}(w)$ and returning the minimum.
[…]
4.3 Analysing the Dynamic Algorithm
In this subsection, we will analyse the dynamic algorithm above. We begin by analysing how precise the estimates $\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v)$ in the relevant data structures are. Then we show that the algorithm actually maintains Invariant 20, thus establishing correctness via Theorem 21. Then we analyse the recourse of the algorithm, before we pick explicit implementations of the data structures and analyse the final running time.
4.3.1 Analysis of Round-Robin
In this section, we will analyse the round-robin scheme applied by the algorithm. In particular, we will show that the algorithm maintains Invariant 20. We will do this by showing that even-though the algorithm sometimes works with (slightly) out-dated information, the information is still precise enough for the algorithm to detect necessary changes and perform them efficiently.
To this end, we first show that the data structures used by the algorithm contain good estimates. Before doing this, we will introduce some notation. For any data structure $\mathcal{D}$, we will let $(\mathcal{D})_{\tau}$ be the state of the data structure after the $(\tau-1)^{th}$ call to ReColour concluded, and just before the $\tau^{th}$ call to ReColour commences. Similarly, for any value $p(x)$ we will let $(\tilde{p}(x))_{\mathcal{D},\tau}$ denote the estimate of this value stored in data-structure $\mathcal{D}$ at time $\tau$, and similarly we will let $(p(x))_{\tau}$ denote the actual value of the parameter at time $\tau$. If the data-structure is clear from the context, we will leave out this argument to ease readability.
Lemma 24.
At all times $\tau$, we have that:
  1. 1.
    For any edge $e=uv$ and any colour $\kappa$, the estimates $(\tilde{S})_{\tau}:=(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau}$ in $T^{s}_{uv}$ satisfy that:
    $|(\tilde{S})_{\tau}-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})|\leq\tfrac{\eta}{16}$
  2. 2.
    For any vertex $u$, any neighbour $v$, and any colour $\kappa$, the estimates
    $(\tilde{E}(u,uv,\kappa))_{\tau}=((\tilde{s}_{c(uv)}(u))+(\tilde{s}_{c(uv)}(v)))-((\tilde{s}_{\kappa}(u))+(\tilde{s}_{\kappa}(v)))_{\tau}$
    stored in $MSD_{\kappa}(u)$ satisfy that:
    $|(E(u,uv,\kappa))_{\tau}-(\tilde{E}(u,uv,\kappa))_{\tau}|\leq\tfrac{\eta}{8}$
    where
    $(E(u,uv,\kappa))_{\tau}=((s_{c(uv)}(u))+(s_{c(uv)}(v)))-((s_{\kappa}(u))+(s_{\kappa}(v)))_{\tau}$
Proof.
We will show item 1)1)1 ) first. Let $\tau^{\prime}<\tau$ be the last time before time $\tau$, where $\tilde{S}$ was updated. We then have that
$(\tilde{S})_{\tau}=(\tilde{S})_{\tau^{\prime}}=(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau^{\prime}}=((s_{\kappa}(u))_{\tau^{\prime}}+(s_{\kappa}(v))_{\tau^{\prime}}$
and so it is sufficient to estimate differences of the form $|s_{\kappa}(w)_{\tau}-s_{\kappa}(w))_{\tau^{\prime}}|$. To this end, we note that due to the round-robin scheme being run every time an edge incident to $w$ changes its colour, we can change $s_{\kappa}(w)$ by at most
$\frac{\Delta_{\max}}{\lceil\tfrac{500\Delta_{\max}}{\eta}\rceil}\leq\tfrac{\eta}{500}$
before it is updated again. As a result, we obtain that $|(s_{\kappa}(w))_{\tau}-(s_{\kappa}(w))_{\tau^{\prime}}|\leq\tfrac{\eta}{32}$, and so the above observations yield:
$|(\tilde{S})_{\tau}-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})|$ $\leq|((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})|$
$\leq|(s_{\kappa}(u))_{\tau}-(s_{\kappa}(u))_{\tau}|+|(s_{\kappa}(v))_{\tau}-(s_{\kappa}(v))_{\tau}|$
$\leq\tfrac{\eta}{16}$
For the second item, first observe that every time $MSD_{\kappa}(u)$ is updated, it is updated with a difference of entrances from some BBST $T^{s}_{ux}$. This means that if $c(uv)=\kappa^{\prime}$, then for all $\kappa$, we actually have that:
$(\tilde{E}(u,uv,\kappa))_{\tau}=(\tilde{s}_{\kappa^{\prime}}(u)+\tilde{s}_{\kappa^{\prime}}(v))_{\tau}-(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau}$
By construction, we also have that:
$(E(u,uv,\kappa))_{\tau}=((s_{\kappa^{\prime}}(u))_{\tau}+(s_{\kappa^{\prime}}(y))_{\tau})-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(y))_{\tau})$
Thus we can apply $1$ to conclude that:
$|\displaystyle|(E(u,uv,\kappa)_{\tau}-(\tilde{E}(u,uv,\kappa))_{\tau}|$$\leq|(\tilde{s}_{\kappa^{\prime}}(u)+\tilde{s}_{\kappa^{\prime}}(v))_{\tau}-((s_{\kappa^{\prime}}(u))_{\tau}+(s_{\kappa^{\prime}}(v))_{\tau})|$
$+|((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})-(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau}|$
$\leq\tfrac{\eta}{16}+\tfrac{\eta}{16}=\tfrac{\eta}{8}$
Next we will argue that this information is accurate enough to maintain Invariant 20, provided that the algorithm terminates. Afterwards, we show that the algorithm always terminates and in fact does so with small amortised recourse.
Lemma 25.
Let $G$ be a dynamic graph, and let $c_{i}$ be the colouring in $G_{i}$ maintained by the algorithm after $i$ updates. Assume that the algorithm terminates after each update and that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then $c_{i}$ satisfies Invariant 20 for all $i$.
Proof.
It suffices to show that after every update, we must have $\Delta s_{\kappa}(w)<\eta$ for all vertices $w$ and all colours $\kappa$. We first note that by Lemma 24, it follows that if $\Delta\tilde{s}_{\kappa}(w)\leq\tfrac{\eta}{2}$, then we have $\Delta s_{\kappa}(w)\leq\Delta\tilde{s}_{\kappa}(w)+\tfrac{\eta}{8}<\eta$.
Hence, if we can show that after the algorithm is done processing any update, we must have $\Delta\tilde{s}_{\kappa}(w)\leq\tfrac{\eta}{2}$ for all vertices $w\in V$ and all colours $\kappa\in[t]$, then we are done. To see this, we apply induction on the number of updates processed. The base case is when no updates have been processed yet in which case what we claim clearly is true. Next assume that the above holds after $i$ updates have been processed, and consider the state of the data structures just after the $(i+1)^{\text{th}}$ update concluded.
[…]
To this end, we note that after a call to Initialize-DS the degree of $w$ is exactly one, and so since $\eta\geq 10$, this call can never cause $\Delta\tilde{s}_{\kappa}(w)>\tfrac{\eta}{2}$.
[…]
Next, we show that the algorithm actually terminates and analyse its recourse.
4.3.2 Analysing the Recourse
In this sub-section, we analyse the recourse of the algorithm and show that the it terminates. We first show that any edge re-coloured by the algorithm results in a significant lowering of the surplus of this edge. This allows us to show that a certain global potential of the graph drops enough to ensure that the amortised number of recolourings become bounded.
To this end, we first argue that any recolouring performed by the algorithm lowers the surplus of an edge significantly:
Lemma 26.
Let $e=uv$ by any edge recoloured by the algorithm during time-step $\tau$. Suppose that $(c(uv))_{\tau-1}=\kappa_{1}$ and that $(c(uv))_{\tau}=\kappa_{2}$ i.e. the colour of the edge is changed from $\kappa_{1}$ to $\kappa_{2}$. Then at time $\tau-1$, we have that
$\delta^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}:=((s_{\kappa_{1}}(u))_{\tau-1}+(s_{\kappa_{1}}(v))_{\tau-1})-((s_{\kappa_{2}}(u))_{\tau-1}+(s_{\kappa_{2}}(v))_{\tau-1})\geq\tfrac{\eta}{4}$
Proof.
Any $\kappa$-coloured edge incident to $u$ or $v$ is only recoloured during time-step $\tau$ if $\Delta\tilde{s}_{\kappa}(u)\geq\tfrac{\eta}{2}$ or $\Delta\tilde{s}_{\kappa}(v)\geq\tfrac{\eta}{2}$. Since $e$ was chosen as an edge maximising the surplus difference, it must be the case that
$\tilde{\delta}^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}$ $:=((\tilde{s}_{\kappa_{1}}(u)+\tilde{s}_{\kappa_{1}}(v))-(\tilde{s}_{\kappa_{2}}(u)+\tilde{s}_{\kappa_{2}}(v)))_{\tau-1}$
$=(\tilde{s}_{\kappa_{1}}(u)+\tilde{s}_{\kappa_{1}}(v))_{\tau-1}-(\tilde{s}_{\kappa_{2}}(u)+\tilde{s}_{\kappa_{2}}(v))_{\tau-1}$
$\geq\tfrac{\eta}{2}$
where the second equality follows from the fact that $MSD_{\kappa}(u)$ always is updated by a difference of surpluses stored in some BBST $T^{s}(uv)$. Hence, it follows that:
$\delta^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}$ $+(\tilde{s}_{\kappa_{1}}(u)+\tilde{s}_{\kappa_{1}}(v))_{\tau-1}-((s_{\kappa_{1}}(u))_{\tau-1}+(s_{\kappa_{1}}(v))_{\tau-1})-((\tilde{s}_{\kappa_{2}}(u)+\tilde{s}_{\kappa_{2}}(v))_{\tau-1}$
$-((s_{\kappa_{2}}(u))_{\tau-1}+(s_{\kappa_{2}}(v))_{\tau-1}))\geq\tilde{\delta}^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}\geq\tfrac{\eta}{2}$
Consequently, we can apply Lemma 24 part 1) in order to observe that:
$\delta^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}$ $\geq\tilde{\delta}^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}-|(\tilde{s}_{\kappa_{1}}(u)+\tilde{s}_{\kappa_{1}}(v))_{\tau-1}-((s_{\kappa_{1}}(u))_{\tau-1}+(s_{\kappa_{1}}(v))_{\tau-1})|-$
$|(\tilde{s}_{\kappa_{2}}(u)+\tilde{s}_{\kappa_{2}}(v))_{\tau-1}-((s_{\kappa_{2}}(u))_{\tau-1}+(s_{\kappa_{2}}(v))_{\tau-1})|$
$\geq\tfrac{\eta}{2}-\tfrac{2\eta}{16}\geq\tfrac{\eta}{4}$
as claimed. ∎
The idea is now to consider the following potential:
$\Phi(G,c)=\sum\limits_{i=1}^{t}\sum\limits_{v\in V(G)}\sum\limits_{j=1}^{s_{i}(v)}j$
As we will show below, Lemma 26 implies that every recolouring performed by the algorithm will lower the potential substantially. In particular, this ensures that the amortised recourse of the dynamic algorithm is low. This is shown in the following Lemma:
Lemma 27.
Let $\tilde{d}=\max_{c}\max_{v}\tilde{s}(v)$ be the maximum surplus of any colouring satisfying Invariant 20 with $10<\varepsilon<$ 1. Then the amortised recourse of the algorithm is upper bounded by $16\frac{\tilde{d}}{\eta}\leq\frac{2^{10}\log n}{\varepsilon^{2}}=O(\tfrac{\log n}{\varepsilon^{2}})$ provided that $\eta=\lfloor\tfrac{\varepsilon^{2}}{128\log n}\rfloor\tfrac{\Delta_{\max}}{t}\geq 16$ and $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$.
In particular, under these assumptions,the algorithm always terminates.
Proof.
We first note that after inserting and colouring any edge $e=uv$, the potential can increase by at most $2(\tilde{d}+1)$, since between insertions the algorithm guarantees that Invariant 20 holds, provided that the algorithm terminates, by Lemma 25. Note that a deletion will always lower the potential.
The second observation is that each recolouring will lower the potential by $\Omega(\eta)$. Indeed, if $e$ is recoloured from having colour $\kappa_{1}$ to having colour $\kappa_{2}$ during time-step $\tau$, then it follows from Lemma 26 that:
$\delta^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}=((s_{\kappa_{1}}(u))_{\tau-1}+(s_{\kappa_{1}}(v))_{\tau-1})-((s_{\kappa_{2}}(u))_{\tau-1}+(s_{\kappa_{2}}(v))_{\tau-1})\geq\tfrac{\eta}{4}$
Notice that if the recolouring performed in step $\tau$ transforms $(G,c)$ to $(G,c^{\prime})$, then we have that:
$\Phi(G,c)-\Phi(G,c^{\prime})=\delta^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}-2\geq\tfrac{\eta}{4}-2\geq\tfrac{\eta}{8}$
Indeed, the above holds since recolouring $e$ reduces $s_{\kappa_{1}}(u)$ and $s_{\kappa_{1}}(v)$ by 1, and hence $\Phi$ by $(s_{\kappa_{1}}(u))_{\tau-1}+(s_{\kappa_{1}}(v))_{\tau-1}$. On the other hand, $s_{\kappa_{2}}(u)$ and $s_{\kappa_{2}}(v)$ are both increased by one, which increases the potential by $(s_{\kappa_{2}}(u)))_{\tau-1}+(s_{\kappa_{2}}(v)))_{\tau-1}+2$.
Since $\Phi$ is always non-negative, since deletions do not lower $\Phi$, and since insertions increase $\Phi$ by at most $2(\tilde{d}+1)$, we find that we can witness at most:
$\frac{r\cdot{}2(\tilde{d}+1)}{\tfrac{\eta}{8}}=16\frac{r\cdot{}\tilde{d}}{\eta}$
recolourings over a sequence of updates containing $r$ insertions. Consequently, the algorithm always terminates and it does so with recourse bounded by $O(\frac{\tilde{d}}{\eta})$. To recover the last part of the Lemma, note that Theorem 21 gives that $\tilde{d}\leq(1+\varepsilon)\tfrac{\Delta_{\max}}{t}$ and that $\eta=\lfloor\tfrac{\varepsilon^{2}}{128\log n}\rfloor\tfrac{\Delta_{\max}}{t}$. ∎
We briefly remark that it is quite important that $\eta\geq 16$. This ensures that any recolouring actually lowers the potential, and that one does not end up in a situation where the algorithm continually recolours the same edges. In the next section, we will analyse the update time as a function of the number of recolourings i.e. the recourse performed by the algorithm. We will then use Lemma 27 to upper bound the total update time.
4.3.3 Analysing the Update Time
We begin by implementing the basic procedures in the following lemma. The proof is straight-forward, but we include a proof for completeness.
Lemma 28.
For vertex $v$, we can support the following operations with deterministic and worst-case guarantees as listed below:
[…]
  • Not counting the recursive call, for all edges $e$, we can perform one call to $\emph{{ReColour}}(e)$ in $O(t^{3}\varepsilon^{-2}\log^{2}m)$ time.
  • We can extract an edge (and the value) maximising $\Delta\tilde{s}_{\kappa}(v)$ in $O(\log m)$ time.
[…]
Proof.
We will only require our BBST’s to support very basic operations, so any standard BBST such as red-black trees [CLRS22] will do.
For the first item, we note that initializing $MSD_{\kappa}(w)$ for any $\kappa\in[t]$ and any $w\in V(G)$ takes constant time since $MSD_{\kappa}$ is initialised as an empty BBST. Initialising the doubly linked-lists over neighbouring edges can also be done trivially in $O(1)$ time for each $\kappa$. Since, we have to do the above for $t$ choices of $\kappa$ and up to $2$ choices of $w$, this part takes $O(t)$ time. We can initialize $T^{s}_{uv}$ in $O(1)$-time and insert the $t$ surplus sums into $T^{s}_{uv}$ in $O(t\log t)$ time. Finally, we can insert the surplus differences into the relevant $MSD_{\cdot{}}(\cdot{})$’s in $O(t\cdot{}\log m)$ time, as $MSD_{\kappa}(w)$ contains $O(m)$ elements for all $\kappa,w$.
For the second item, we note that each update to $T^{s}_{uv}$ takes $O(\log t)$ time, since we can compute $s_{\kappa}(u)+s_{\kappa}(v)$ in constant time. Each update to $MSD_{\kappa}(v)$ can also be performed in $O(\log n)$ time. Indeed, we can compute $(s_{c(uv)}(u)+s_{c(uv)}(v))-(s_{\kappa}(u)+s_{\kappa}(v))$ in $O(1)$ time, and insert it in $O(\log m)$ time. Similarly, by storing a pointer to the element at $w$, we can also remove the old entrance in $O(\log m)$ time. All of the above has to be performed $t$ times, leaving us with the stated update time.
[…]
For the fifth item, we note that by storing a pointer from each entrance in $MSD_{\kappa}(v)$ to the edge responsible for the entrance, we can, given access to an entrance, locate the corresponding edge in $O(1)$ time. As such, we only need to compute the maximum entrance in $MSD_{\kappa}(v)$ which can be done in $O(\log m)$ time.
For the sixth and final item, we note that extracting the maximum entrance in a BBST can be performed in $O(\log n)$ time by a standard search always going to the largest child. ∎
Everytime we call ReColour on some edge, we perform a recolouring. Hence, by Lemma 27, after $i$ insertions, we have performed at most $16\cdot{}i\cdot{}\frac{\tilde{d}}{\eta}$ calls to ReColour. Thus, in an amortised sense, the number of calls to ReColour is $O(\frac{\tilde{d}}{\eta})$ per update. If we ignore the recursive calls, then by Lemma 28 the time spent for each call is dominated by the call to Round-Robin, and so each call to Recolour takes only $O(t^{3}\varepsilon^{-2}\log^{2}m)$ time. Hence, we have shown the following bound on the amortised running time of insert and delete.
Lemma 29.
The amortised running time of insert and delete is upper-bounded by $O(t^{3}\varepsilon^{-4}\log^{3}m)$.
Combining Lemma 29, Lemma 27, and Lemma 25 now gives Theorem 3.
5 Maintaining a hierarchy of $t$-splitters
In this section, we will show how to build a hierarchy of dynamic $t$-splitters in order to split the degree all the way down to $O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$ while retaining an update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$. Thus we will show Theorem 4, which we restate below:
Theorem 30 (Identical to Theorem 4).
Let $G$ be a dynamic graph, and let $01>\varepsilon>$ 0 and an upper bound $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ on the maximum degree of $G$ through the entire update sequence be given. Set $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$ and $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Then there exists a dynamic algorithm maintaining parameters $h,i,j$ with $h=i+j\leq 4\sqrt{\log n}$ and a set of graphs $\mathcal{G}_{h}$ satisfying that:
  1. 1.
    $|\mathcal{G}_{i+j}|\leq t_{1}^{i}t_{2}^{j}$
  2. 2.
    $\hat{\Delta}_{i+j}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
The algorithm has amortised recourse and update time both in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$.
In this section, we will again assume we are given an upper bound $\Delta_{\max}$ on $\Delta$. We first give a formal description of the hierarchy and an algorithm that maintains it.
5.1 The hierarchy
We first describe the hierarchy in words. The hierarchy will consist of two types of $t$-splitters. The first type of $t$-splitter, we will use, has $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$, and the second type has $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Given any graph $G$, any $\Delta_{\max}$, and any $10<\varepsilon<$ 1, we will select some $h$. The hierarchy then consists of a sequence of sets of graphs $\{\mathcal{G}_{i}\}_{i=0}^{h}$ and a sequence of values $\{\hat{\Delta}_{i}\}_{i=0}^{h}$ such that for all $0\leq i\leq h$ and all $H\in\mathcal{G}_{i}$, we have that $\max_{v\in V(H)}d(v)\leq\hat{\Delta}_{i}$ i.e. the sequence of $\hat{\Delta}_{i}$ provides a sequence of upper bounds on the max degrees.
We note already here that since the algorithm from Theorem 3 is not history independent and so ’running’ it on a static graph is ill-defined. We will ignore this in the following informal description, and then provide a proper definition afterwards. We will initialise $\mathcal{G}_{0}=\{G\}$ and $\hat{\Delta}_{0}=\Delta_{\max}$, and proceed as follows: given $\mathcal{G}_{i}$ and $\hat{\Delta}_{i}$, we let $\mathcal{G}_{i+1}$ be the set of graphs obtained in the following manner. If $\Delta_{\max}\geq 2^{10i\sqrt{\log n}}$, we will run the algorithm from Theorem 3 with the parameter $\eta=\tfrac{\varepsilon^{\prime 2}}{\log n}$ and $t=t_{1}$ on all $H\in\mathcal{G}_{i}$. For each $H$, every colour-class $\kappa$ induces a subgraph $H_{\kappa}$ of $H$. For each choice of $H$ and $\kappa$, we include $H_{\kappa}$ in $\mathcal{G}_{i+1}$ after removing all degree 00 vertices. Finally, by Theorem 3, we can choose $\hat{\Delta}_{i+1}=(1+\eta)\tfrac{\hat{\Delta}_{i}}{t}$. Otherwise we set $t=t_{2}$ and repeat the above construction. We will pick $h$ such that the construction terminates with $\hat{\Delta}_{h}=O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$, but still large enough for any application of Theorem 3 to go through.
The idea is that the edge sets of the graphs in $\mathcal{G}_{h}$ form a partition of the edges of $G$ such that each partition induces a low-degree subgraph of $G$. Since the $t$-splitters used are fully-dynamic, we can maintain the above hierarchy dynamically as well (see Algorithms 7 8 9). To do so, we simply build the $t$-splitters as specified, and whenever an edge is inserted, deleted or recoloured, we remove it from the graphs it previously belonged to (if any) and instead insert it into the graphs specified by its new colour (if any). The order of operations does not matter too much, as long as we guarantee that any $t$-splitter algorithm that is called is allowed to finish and that the $\hat{\Delta}_{i}$’s remain valid upper-bounds through-out the update sequence. In Algorithms 7 8 9 we will explicitly fix the order of operations.
Next, we introduce some basic notation in order to make the above description formal. First of all, we define a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy. This is any hierarchy obtained by splitting via any of $t$-splitter satisifying Invariant 20, and not necessarily only via the dynamic $t$-splitter from Theorem 3. This deals with the technicality that the data structure from Theorem 3 might not be history-independent, and so we cannot dynamically hope to maintain a very strictly defined hierarchy.
Definition 31.
Given $G,\Delta_{\max}$, and $\varepsilon$. A $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy of depth $h$ is a sequence of sets of graphs $\{\mathcal{G}_{i}\}_{i=0}^{h}$, a sequence of sets of colourings $\{\mathcal{C}_{i}\}_{i=0}^{h}$, and two sequences of integers $\{\hat{\Delta}_{i}\}_{i=0}^{h}$ and $\{t(i)\}_{i=0}^{h-1}$ such that:
  • $\mathcal{G}_{0}=\{G\}$ and $\hat{\Delta}_{0}=\Delta_{\max}$.
  • For all $i$: if $\Delta_{\max}\geq\lfloor 2^{10\sqrt{\log n}}\rfloor\lfloor 2^{\sqrt{\log n}}\rfloor^{i-1}$ then $t(i)=t_{1}$, else $t(i)=t_{2}$.
  • For each $H\in\mathcal{G}_{i}$ there is a corresponding $t(i)$-splitting colouring $c_{H}\in\mathcal{C}_{i}$.
  • $\mathcal{G}_{i+1}:=\{\hat{H}:\exists H\in\mathcal{G}_{i},\kappa\in[t^{\prime}]\text{ with }H[\kappa]=\hat{H}\text{ under }c_{H}\}$.
  • For all $i$: $\hat{\Delta}_{i+1}\leq(1+\tfrac{\varepsilon}{128\log n})\tfrac{\hat{\Delta}_{i}}{t^{\prime}}$
  • For all $i$, all $H\in\mathcal{G}_{i}$, and all $v\in V(H)$: $d_{H}(v)\leq\hat{\Delta}_{i}$
Remark 32.
We remark that the choice of the value $2^{(10+(i-1))\log n}$ is not extremely precise. One could have equally well chosen $2^{(9+(i-1))\log n}$ or $20\cdot{}2^{(10+(i-1))\log n}$. One then need only change the constant in Algorithms 7 9 8, in Definition 31, and in the subsequent proofs mutatis mutandi. In particular, one can always discretize the values and slightly over- or under-estimate all parameters like $\Delta_{\max},\varepsilon$ or $2^{(10+(i-1))\log n}$ to ensure that one does not need access to real RAM in order to do fast comparisons.
In particular, any hierarchy that uses a valid instantiation of Theorem 3 for every $t$-splitter yields a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy; regardless of the insertion order. We will also observe that for all $i$, the edge sets of the graphs in $\mathcal{G}_{i}$ form a valid partition of $E(G)$. We will let a vector $\chi(e)$ describe the hierarchy-profile of an edge:
Definition 33.
For any edge $e$ and any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy, we let the hierarchy-profile of $e$ be the vector $\chi(e)\in\mathbb{Z}^{h}$ such that if $e\in E(H)$ for $H\in\mathcal{G}_{i}$, then $\chi(e)_{i}=c_{H}(e)$.
For any $i\leq h$ and any vector $v\in\mathbb{Z}^{h}$, we let the $i^{\text{th}}$ initial segment of $v$ be the vector $v^{(i)}\in\mathbb{Z}^{i}$ whose first $i$ coordinates agree with the first $i$ coordinates of $v$. For any vector $v\in\mathbb{Z}^{i}$ and constant $k\in\mathbb{Z}$, we also let $[v^{(i)},k]$ be the vector in $\mathbb{Z}^{i+1}$ with $i^{th}$ initial segment equal to $v$ and $[v,k]_{i+1}=k$ We can index the graphs in the hierarchy by providing an index set $J$ containing every initial segment of the hierarchy-profile of every edge $e\in G$. Here, we let $G_{\emptyset}=G$, and we let $G_{j}$ for any $j\in J$ be the graph containing exactly the edges $e$ for which $j$ is an initial segment of $\chi(e)$. We let $|j|:=dim(j)$ denote the dimension of the vector $j$. Also $|j|$ is the level of $G_{j}$.
5.2 An algorithm for maintaining a hierarchy
The algorithm for dynamically maintaining the hierarchy is quite straight-forward. First of all, one initializes values $t(i)$ and $\hat{\Delta}_{i}$ so that they conform to the constraints of Definition 31 (see Algorithm 7). This only has to be done once.
The hierarchy is then maintained by repeatedly applying Theorem 3 to further split graphs into ones with lower degree. First one performs the update on level 00, which then propagates the update along with any extra recolourings to the next level of the hierarchy. One has to be a bit careful with determining the order of deletions and insertions, in order to ensure that the $\hat{\Delta}_{i}$’s remain valid upper-bounds through-out the update sequence. Therefore, the algorithm first processes an update in the top level, deletes all edges that were recoloured from the old graphs they were put into, and then inserts any new edges into the further recursive structures. This ensures that $\hat{\Delta}_{i}$’s remain valid upper-bounds. Finally, we note that the hierarchy consists of many graphs, which can either be constructed in a pre-processing phase or simply allocated online by only inserting vertices into a graph once the vertex actually has degree at least one.
Next, we give pseudo-code for the above dynamic algorithm. If an edge $e$ has just been recoloured, we let $c(e)_{\text{old}}$ respectively $\chi(xy)_{\text{old}}$ denote the colours of $e$ just before the recolouring. The hierarchy is maintained via the following set of recursive procedures. The first procedure shows how to initialize the parameters needed for the following procedures. We note already now that to actually perform an insertion or a deletion, one must call the respective recursive procedure with the correct parameter setting (see Remark 34):
  $i_{1}=0$
  $i_{2}=0$
  $\hat{\Delta}_{0}=\Delta_{\max}$
  $\mathcal{G}_{0}=\{G\}$
  $\eta=\lfloor\tfrac{\varepsilon}{128\log n}\rfloor$
  while $\hat{\Delta}_{i_{1}+i_{2}}\geq\lceil\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}\rceil$ do
     if $\tfrac{\Delta_{\max}}{\lfloor 2^{\sqrt{\log n}}\rfloor^{i_{1}-1}}\geq\lfloor 2^{10\sqrt{\log n}}\rfloor$ then
        $t(i_{1}+i_{2})=t_{1}$
        $i_{1}=i_{1}+1$
        $\hat{\Delta}_{i_{1}+i_{2}}=(1+\eta)^{i_{1}+i_{2}}\tfrac{\Delta_{\max}}{t_{1}^{i_{1}}t_{2}^{i_{2}}}$
     else
        $t(i_{1}+i_{2})=t_{2}$
        $i_{2}=i_{2}+1$
        $\hat{\Delta}_{i_{1}+i_{2}}=(1+\eta)^{i_{1}+i_{2}}\tfrac{\Delta_{\max}}{t_{1}^{i_{1}}t_{2}^{i_{2}}}$
  $h=i_{1}+i_{2}$
Algorithm 7 Initialize-Parameters$(\Delta_{\max},\varepsilon)$
  insert(uv) into $G_{j}$ using Theorem 3 with $(\Delta_{\max},t,\varepsilon^{\prime})=(\hat{\Delta}_{|j|},t(|j|),\eta)$
  for all recoloured edges $xy\neq uv$ in $G_{j}$ do
     for $k=h$ to $h-|j|$ do
        Delete-Hierarchy$(xy,\chi(xy)^{(k)}_{\text{old}},\chi(xy)_{\text{old}})$
     Insert-Hierarchy$(xy,[j,c_{G_{j}}(xy)])$
  Insert-Hierarchy$(uv,[j,c_{G_{j}}(uv)])$
Algorithm 8 Insert-Hierarchy$(uv,j)$
  delete(uv) from $G_{j}$ using Theorem 3 with $(\Delta_{\max},t,\varepsilon^{\prime})=(\hat{\Delta}_{|j|},t(|j|),\eta)$
  if $|v|\geq|j|+1$ then
     Delete-Hierarchy$(uv,v^{(|j|+1)},v)$
  for all recoloured edges $xy\neq uv$ in $G_{j}$ do
     for $k=h$ to $h-|j|$ do
        Delete-Hierarchy$(xy,\chi(xy)^{(k)}_{\text{old}},\chi(xy)_{\text{old}})$
     Insert-Hierarchy$(xy,[j,c_{G_{j}}(xy)])$
Algorithm 9 Delete-Hierarchy$(uv,j,v)$
Remark 34.
[…]
5.3 Correctness
In this section, we show that the algorithms from above actually maintain a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with parameters that give the guarantees stated in Theorem 4. We first show that the algorithm maintains a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy:
Lemma 35.
Algorithms 7 8 9 implement a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy.
Proof.
[…]
The hierarchy is also in accordance with the third, fifth, and sixth items. Indeed, if at all times the conditions of Theorem 3 is met, then it follows by induction and the theorem itself that the computed colourings are $t(i)$-splitting for all $i$, and furthermore satisfy the fifth and sixth items.
To check that this is the case, we check that the upper bounds on the degrees $\hat{\Delta}_{i}$ remain valid at all times during an update. This holds since the algorithms immediately delete all further copies of an edge upon a deletion or recolouring. Indeed, suppose the graph $G_{v}$ is a graph with $|v|$ minimum that invalidates the upper-bound $\hat{\Delta}_{|v|}$. Then it must be because the colouring on $G_{|v|^{(|v|-1)}}$ is not a $t(|v|-1)$-splitter, or because $G_{v}$ at some point is not a subgraph of $G_{|v|^{(|v|-1)}}[v_{|v|}]$. The first case cannot happen, since the $t(|v|-1)$-splitter on $G_{|v|^{(|v|-1)}}$ always has all requirements satisfied by assumption. The second case cannot happen, since any time an edge coloured $v_{|v|}$ is deleted or recoloured in $G_{|v|^{(|v|-1)}}$, it is immediately removed from $G_{v}$. Hence, any time that $G_{v}$ is subjected to an insertion, it is a subgraph of $G_{|v|^{v}}[v_{|v|}]$ and thus satisfies the degree-bound.
Since we construct $\mathcal{G}_{i+1}$ by applying Theorem 3 to every graph in $\mathcal{G}_{i}$, and since Theorem 3 returns a valid $t(i)$-splitter, we thus also recover Item $4$. ∎
We will use the following standard estimate repeatedly:
Claim 36.
For $10<\varepsilon<$ 1 and $0<\eta\leq\tfrac{\varepsilon}{128\log n}$, we have that $(1+\eta)^{4\log n}\leq(1+\tfrac{\varepsilon}{16})$
Proof.
It follows from the well-known fact that $1+r\leq e^{r}$ for all $r\in\mathbb{R}$ that $(1+\tfrac{x}{q})^{q}\leq e^{x}$ for all reals $x,q\neq 0$. Indeed, we can write;
$(1+\eta)^{4\log n}\leq(1+\tfrac{\varepsilon}{128\log n})^{4\log n}\leq e^{\tfrac{\varepsilon}{32}}$
Furthermore, for $r>-1$ it is well-known that $\tfrac{r}{1+r}\leq\log(1+r)$ and so we have
$(1+\eta)^{4\log n}\leq e^{\tfrac{\varepsilon}{16(1+\tfrac{\varepsilon}{16})}}\leq(1+\tfrac{\varepsilon}{16})$
as claimed. ∎
Next, we show that:
Lemma 37.
For any graph $G$ with at least one edge and any upper bound $\Delta_{\max}\leq n$, we have that any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy of depth $h$ must have $h\leq 4\sqrt{\log n}$.
Proof.
Observe that we can perform at most $i\leq 2\sqrt{\log n}$ splits with $t=t_{1}$. Indeed, after $i$ splits, we have $\tfrac{\Delta_{\max}}{t_{1}^{i}}\leq\tfrac{n}{t_{1}^{i}}$, so setting $i>2\sqrt{\log n}$ forces $\tfrac{\Delta_{\max}}{t_{1}^{i-1}}<\lfloor 2^{10\sqrt{\log n}}\rfloor$ and so $\Delta_{\max}<\lfloor 2^{10\sqrt{\log n}}\rfloor\cdot{}\lfloor 2^{\sqrt{\log n}}\rfloor^{i-1}$. This means that $t(j)\neq t_{1}$ for all $j\geq i$.
A similar argument shows that we perform at most $2\sqrt{\log n}$ splits with $t=t_{2}$ before $\tfrac{2^{10\sqrt{\log n}}}{t_{2}^{i-1}}$ goes below $1/2$. By Lemma 36, we have $(1+\eta)^{4\sqrt{\log n}}<2$, and so $\hat{\Delta}_{4\sqrt{\log n}}<1$, contradicting the fact that $\hat{\Delta}_{2\sqrt{\log n}}$ must upper bound the maximum degree of all graphs in $\mathcal{G}_{4\sqrt{\log n}}$, one of which must have degree at least one, since $\mathcal{G}_{2\sqrt{\log n}}$ partitions $G$. ∎
Finally, we show that any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ satisfies the conditions of Theorem 4. We note that if $\Delta_{\max}<\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$, we do not need to apply the hierarchy, as we can simply apply the techniques develop in Section 6 directly on the graph.
Lemma 38.
Given any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$. Let $i$ be the smallest integer $i$ for which $\tfrac{\Delta_{\max}}{t_{1}^{i}}<\lfloor 2^{10\sqrt{\log n}}\rfloor$. Similarly, let $j$ be the smallest integer $j$ for which $\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}<\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ Then, assuming the hierarchy has depth $i+j$, we have:
  1. 1.
    $|\mathcal{G}_{i+j}|\leq t_{1}^{i}t_{2}^{j}$
  2. 2.
    $\hat{\Delta}_{i+j}=(1+\eta)^{i+j}\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
Proof.
We show the first item first. Observe that we perform $i$ splits into at most $t_{1}$ graphs first, meaning that $|\mathcal{G}_{i}|\leq t_{1}^{i}$ by an easy induction on $i$, since $|\mathcal{G}_{0}|=1$. A similar induction on $j$ yields the result, since this time each split is into $t_{2}$ graphs. The proof of the first part of the second item is a completely synchronous induction on first $i$ and then $j$ using item 5 of Definition 31.
The last in-equality of item $2$ follows by observing that $i+j\leq 4\sqrt{\log n}$ by Lemma 37 and then applying Claim 36. ∎
Lemma 35 and Lemma 38 together establish correctness of the Algorithm. In order to prove Theorem 4, we need now only analyse the running time and the recourse of the hierarchy. We do so in the next sub-section.
5.4 Analysis
The following Lemma analyses the recourse the algorithm for maintaining the hierarchy:
Lemma 39.
Starting from an empty graph, a sequence of $\sigma$ insertion or deletions to graphs at level $i$ triggers at most $\sigma\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i}$ insertions or deletions to graphs at level $j\geq i$.
Proof.
The proof is by strong induction on $j-i$. The base case where $j-i=0$ clearly holds. We proceed to the induction step. Assume that for $j-i=0$, $j-i=1$, …, $j-1=k$ the lemma holds. Now assume $j-i=k+1$. First of all, we note that any insertion or deletion at a lower level might trigger a similar insertion or deletion at all higher levels. By the induction hypothesis, the number of such insertion or deletions, $U_{1}$, at level $j$ is bounded as follows:
$U_{1}\leq\sum\limits_{\ell=i}^{j-1}\sigma\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{\ell-i}\leq\tfrac{\sigma}{2}\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i}$
since $\sum_{i=1}^{y}x^{i}\leq\tfrac{x^{y+1}}{x-1}$ for $x\geq 2$. Note that these insertions and deletions account not only for direct insertions or deletions into $G$, but also for insertions and deletions resulting from re-colourings at some level $\ell<j-1$. Indeed, these re-colourings trigger insertions and deletions at higher levels, whose contributions are already accounted for in the quantity above.
Thus it only remains to account for the number of insertions or deletions, $U_{2}$, caused by re-colourings at level $j-1$. Observe that by Lemma 27 it follows that over any sequence of $\sigma^{\prime}$ insertions or deletions to a graph at level $j-1$, at most $\sigma^{\prime}\cdot{}\tfrac{2^{10}\log n}{\varepsilon^{2}}$ recolourings are performed at level $j-1$. Each recolouring corresponds to one insertion and one deletion on the next level, and so we find that at most
$U_{2}\leq 2\cdot{}\sigma\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-1-i}\cdot{}\tfrac{2^{10}\log n}{\varepsilon^{2}}\leq\tfrac{\sigma}{2}\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i}$
insertions or deletions are performed at level $j$ as a result of re-colourings at level $j-1$. Summing $U_{1}$ and $U_{2}$ finishes the induction. ∎
The following lemma is then immediate:
Lemma 40.
The amortised recourse of an insertion or deletion to $G$ is in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$. Furthermore, the amortised update time is also in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$.
Proof.
Starting from any empty graph undergoing a sequence of $\sigma$ updates Lemma 39 and Lemma 37 together imply that at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ insertion and deletions are performed at level $\ell$ for any $\ell\leq h$. Since the hierarchy maintained has depth $h\leq 4\sqrt{\log n}$ by Lemma 37, an insertion or deletion triggers at most
$\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$
re-colourings, insertion or deletions to some level of the hierarchy. Hence, over the sequence of $\sigma$ updates the recourse is bounded by $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$.
To recover the second part of the lemma, we know by above that, starting from an empty graph, any sequence of $\sigma$ operations trigger at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ insertion and deletions. Hence, for the usual index set $I$, we know that if $G_{v}$ experience $\sigma_{v}$ insertion or deletions, then
$\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}=\sum_{v\in I}\sigma_{v}$
By Theorem 3, we can process a sequence of $\sigma_{v}$ insertions or deletion to a graph of level $|v|$ in $c\sigma_{v}\log(m)t_{1}^{3}$ update time for some constant $c$. As a result, the total update time is
$\sum_{v\in I}c\sigma_{v}\log(m)t_{1}^{3}=c\log(m)t_{1}^{3}\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}=\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$
and so the amortised update time is as claimed. ∎
[…]
6 Deterministic Dynamic Edge-Colouring of Low-Degree Graphs
In this section, we show how to deterministically maintain a $(\Delta_{\max}+1)$-edge-colouring in a dynamic graph with maximum degree $\Delta\leq\Delta_{\max}$. In Section 7.2, we will show how to adjust the algorithm to the case where no upper bound on $\Delta$ is needed, at the cost of the algorithm using $(1+\varepsilon)\Delta$ colours instead. We will follow the approach sketched in the informal overview. We begin by showing some structural corollaries of previous work:
6.1 Structural Corollaries
As mentioned earlier, a key observation is that a particular bichromatic path cannot be reached from too many different starting points. Similar observation has been used in the literature before [Ber22, BD23, Chr23] to analyse algorithms. Here, we will apply the fact more explicitly and actually update data structures at every point that may reach a recoloured edge via a short and non-overlapping multi-step process with few steps. We will use the following Lemma due to Christiansen [Chr23]. Note that originally the lemma is stated using the maximum degree $\Delta$, but the Lemma carries over ad-verbatim if one replaces $\Delta$ by an upper bound $\Delta_{\max}$ on $\Delta$:
Lemma 41 (Lemma 26 in full-version of [Chr23]).
Let $H\subset G$ be a subgraph of $G$ and let $c,c^{\prime}$ be two (partial) $(\Delta_{\max}+1)$-edge-colourings of $G$ differing only on $H$. Then, under both $c$ and $c^{\prime}$, the same at most $2(\Delta_{\max}+1)^{3}\cdot{}|H|$ points in $G-N^{1}(H)$ can reach $H$ via a 1-step process.
We have the following Corollary 42 below:
Corollary 42.
Let $G$ be a graph with maximum degree $\Delta\leq\Delta_{\max}$ endowed with a proper (possibly partial) $(\Delta_{\max}+1)$ edge-colouring $c$. Let $v\in V(G)$ be any vertex. Then $v$ can be reached by at most $3^{i}(\Delta_{\max}+1)^{3i}$ non-overlapping $i$-step processes.
Furthermore, if $C$ is a semi-consistent $j$-step process with step-length $\leq\ell$, then $N^{2}(C)$ can be reached by at most $3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}j\cdot{}\ell$ non-overlapping $i$-step processes.
Proof.
We show both claims via induction on $i$. We begin with the first claim. To this end, fix an arbitrary vertex $v$. Then $v$ can be reached by at most $2(\Delta_{\max}+1)^{3}$ points in $G-N^{1}(v)$ via a bichromatic path by Lemma 41. Any point in $N^{1}(v)$ might also reach $v$ by simply hopping to a neighbour, and so we conclude that in total at most $2(\Delta_{\max}+1)^{3}+(\Delta_{\max}+1)\leq 3(\Delta_{\max}+1)^{3}$ points are able to reach $v$ by a 1-step process. This concludes the base case. Now let $w$ be some point that can reach $v$ via $(i-1)$-steps of the above process. Then by applying the same argument as above, we find that at most $3(\Delta_{\max}+1)^{3}$ points can be the basis of a $1$-step process that can reach $w$. By applying induction we find that there are at most $3(\Delta_{\max}+1)^{3(i-1)}$ choices for $w$, and so since any $i$-step process reaching $v$ has to go through $w$, we find that there are at most
$3(\Delta_{\max}+1)^{3}\cdot{}3^{i-1}(\Delta_{\max}+1)^{3(i-1)}=3^{i}(\Delta_{\max}+1)^{3i}$
such points, which is what we wanted to show.
The second claim follows a similar argument to above. We will let $H=N^{2}(C)$ be the one-hop neighbourhood of the $j$-step process, and focus on the first vertex in $H$ reached by some process. Then we can apply Lemma 41 and the same argument as above to find that at most $3(\Delta_{\max}+1)^{3}|H|$ points can reach $H$ via a $1$-step process. The induction step is similar to above, except for the fact that we now have at most $3^{i-1}(\Delta_{\max}+1)^{3(i-1)}|H|$ choices for $w$, and so this time we find that there at most
$3(\Delta_{\max}+1)^{3}\cdot{}3^{i-1}(\Delta_{\max}+1)^{3(i-1)}|H|=3^{i}(\Delta_{\max}+1)^{3i}|H|=3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}j\cdot{}\ell$
of the desired points, which is what we wanted to show. ∎
Next we will need the following notation. Note that the definitions deviate slightly from those presented in [Chr23], as we now require the Vizing chains to be consistent according to Definition 14.
Definition 43 ([Chr23]).
Given a proper partial colouring and an uncoloured edge $e$, let $v\in R_{i}(e,\ell)$ denote the set of vertices such that there exists a consistent and non-overlapping $i$-step Vizing chain from $e$ ending at $v$ such that every Vizing chain in the $i$-step Vizing chain has length at most $\ell$.
Definition 44 ([Chr23]).
Given a proper partial colouring and an uncoloured edge $e$, let $v\in N_{i}(e,\ell)\subset R_{i}(e,\ell)$ if there exists a consistent and non-overlapping $i$-step chain from $e$ ending at $v$, and, furthermore, the multi-step Vizing chain may be extended to a non-overlapping $(i+1)$-Vizing chain by concatenating a Vizing chain of length at most $\ell$ at $v$.
In order to make it easier to maintain the data structures used to find short augmenting multi-step Vizing chains, we have imposed some further restrictions on how the chains are allowed to be extended. Most of these restrictions are already implicitly applied in for instance Christiansen’s analysis of multi-step Vizing chains [Chr23]. While this analysis allow for Vizing chains not adhering to the restrictions, we have imposed, all arguments essentially impose most of these restrictions on the chains anyway, and so the arguments go through with the above restrictions added almost ad verbatim. Hence, the following lemma is implicit in [Chr23]:
Lemma 45 (Implicit in Lemma 27 in full-version of [Chr23]).
Suppose we are given a proper partial colouring and an uncoloured edge $e$ in a graph of maximum degree $\Delta\leq\Delta_{\max}$, and let $\ell\geq 80(\Delta_{\max}+1)^{7}$. Suppose that for all $11\leq j\leq i-$ 1 it holds that no vertex in $N_{j}(e,\ell)$ can be extended to an augmenting and non-overlapping $(j+1)$-step Vizing chain via a Vizing chain of length at most $\ell$. Suppose furthermore that the conditions below hold for all $11\leq j\leq i-$ 1:
  1. 1.
    $|N_{j}(e,\ell)|\geq\frac{\ell}{8\Delta_{\max}^{3}}\sum\limits_{k=1}^{j-1}|N_{k}(e,\ell)|$
  2. 2.
    $|N_{j}(e,\ell)|\geq|R_{j}(e,\ell)|/2$.
Then condition 1)1)1 ) and 2)2)2 ) also hold for $j=i$.
The lemma allows us to deduce the existence of short augmenting and consistent multi-step Vizing chains taking a sub-$\log n$ number of steps. This follows from arguments very similar to those from [Chr23], just with a different choice of parameters:
Corollary 46.
Let $G$ be any simple graph with maximum degree $\Delta\leq\Delta_{\max}$. Let $c$ be any proper partial $(\Delta+1)$-edge-colouring of $G$. Then for any uncoloured edge $e$, under the colouring $c$, there is an augmenting, consistent and non-overlapping Vizing chain with at most $\sqrt{{\log n}}$ steps and total length at most $80(\Delta_{\max}+1)^{7}\cdot{}2^{\sqrt{\log n}}\sqrt{\log n}$.
Proof.
We want to apply Lemma 45 recursively until the vertex set is exhausted, and the assumption of the Lemma no longer holds. This will then imply the existence of the object we seek. To this end, we will fix $\ell=80(\Delta_{\max}+1)^{7}\cdot{}2^{\sqrt{\log n}}$ and fix an arbitrary uncoloured edge $e$.
We first construct an augmenting Vizing chain on $e$ as specified in Section 3, making sure that it is consistent with Definition 14. If it has length $\leq\ell$, we are done, so we may assume this is not the case. Next we consider the action of truncating the Vizing chain at any of the $\ell$ first vertices, and extending it via Lemma 13 in a manner consistent with Definition 14.
At most $2\Delta_{\max}$ of these extensions will overlap the edges of the path of the first Vizing chain – namely if the bichromatic path they extend through coincides with the bichromatic path of the Vizing chain build on $e$. At most $(\Delta_{\max}+1)^{2}$ bichromatic paths can go through a vertex, and therefore at most $2(\Delta_{\max}+1)^{4}$ of the extensions will intersect the one-hop neighbourhood of the center of the fan situated at $e$. Hence we find that $|R_{1}(e,\ell)|=\ell$ by construction and that $|N_{1}(e,\ell)|\geq\ell-2\Delta_{\max}-2(\Delta_{\max}+1)^{4}\geq\ell/2$. Hence, we may apply Lemma 45 until we find a short augmenting and non-overlapping Vizing chain. In particular, it follows by induction that on $j$ that
$|N_{j}|\geq\mathopen{}\left(\frac{\ell}{8\Delta_{\max}^{3}}\right)\mathclose{}^{j}$
provided that for every $i\leq j$, we have no augmenting and consistent $i$-step Vizing chain. In particular, since $N_{j}\subset V(G)$ for all $j$, we conclude that
$\mathopen{}\left(\frac{\ell}{8\Delta_{\max}^{3}}\right)\mathclose{}^{j}\leq|V(G)|\leq n$
and consequently we have:
$i\leq\log_{\frac{\ell}{8\Delta_{\max}^{3}}}n\leq\frac{\log n}{\log\frac{\ell}{8\Delta_{\max}^{3}}}\leq\frac{\log n}{\sqrt{\log n}}=\sqrt{\log n}$
In particular, this means that there must exist some augmenting and consistent Vizing chain with at most $\sqrt{\log n}$ steps and length at most $\ell\cdot{}\sqrt{\log n}=80(\Delta_{\max}+1)^{7}\cdot{}2^{\sqrt{\log n}}\sqrt{\log n}$
At the same time, it follows from Corollary 46 and Lemma 17, that at most $2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$ of the $i$-step consistent and non-overlapping Vizing chains can reach the Vizing chain. The choice of $\ell=80(\Delta_{\max}+1)^{7}\cdot{}2^{\sqrt{\log n}}$ allows us to set $i\leq\sqrt{\log n}$ while ensuring the existence of augmenting, consistent and non-overlapping $i$-step Vizing chains by Corollary 46. This guarantees the Vizing chain has length at most $2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$.
At the same time, it follows from Corollary 42 and Lemma 17, that at most $2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$ of non-overlapping $i$-step processes can reach the Vizing chain. This allows us to exhaustively locate all of these chains if necessary. The combination of these properties is what allows us to build a dynamic data structure for implicitly maintaining a shortest augmenting and consistent Vizing chain for any edge that might be inserted into the dynamic graph. In the next section we will describe the data structures needed to do so.
6.2 Contents of the data structures
We will describe the data structures in terms of the following graph:
Definition 47.
Let $G$ be a graph with maximum degree $\Delta\leq\Delta_{\max}$, let $c$ be a proper (partial) $(\Delta_{\max}+1)$-edge-colouring of $G$, and let $\ell$ be a parameter. The bichromatic skeleton of $G$ is the directed graph $\mathcal{B}(G,c,\ell)$ with:
  • Vertex-set $V(\mathcal{B}(G,c,\ell))=X_{1}\cup X_{2}$, where every maximal bichromatic path $P$ of $G$ under $c$ is represented by a node in $X_{1}$ and every node in $V(G)$ is represented by a node in $X_{2}$. Note that we allow $P$ to consist of an isolated vertex.
  • Edge-set $E(\mathcal{B}(G,c,\ell))=Y_{1}\cup Y_{2}\cup Y_{3}$. Here:
  • $e=(x_{2}x_{1})\in Y_{1}$ with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ if $x_{2}$ is an endpoint of the bichromatic path represented by $x_{1}$.
  • $e=(xy)\in Y_{2}$ with $x,y\in X_{2}$ if the nodes in $G$ represented by $x,y$ in $\mathcal{B}(G,c)$ are neighbours via a coloured edge in $G$.
  • $e=(x_{1}x_{2})\in Y_{3}$ with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ if $x_{2}$ belongs to the vertex set of the bichromatic path represented by $x_{1}$, and, furthermore, the distance in the bichromatic path, represented by $x_{1}$, from $x_{2}$ to an endpoint of the bichromatic path is at most $\ell$
If $c$ and $\ell$ is clear from the context, we will often leave the arguments out and simply write $\mathcal{B}(G)$. Let $\mathcal{P}(G,c,\ell)$ or simply $\mathcal{P}$, if the arguments are clear from the context, be the set containing all bichromatic paths $P$ that are contained in some maximal bichromatic path $P^{\prime}$ of $G$ under $c$ such that $P$ has length at most $\ell$ and such that $P\cap P^{\prime}$ contains an endpoint of $P^{\prime}$. Then there is a canonical map $r:\mathcal{P}\cup V(G)\rightarrow V(\mathcal{B}(G))$ that maps a vertex in $V(G)$ to its representative in $X_{2}$, and similarly a bichromatic path $P$ to the representative in $X_{1}$ of the unique maximal bichromatic path $P^{\prime}$ containing $P$. Below we state some very basic and straight-forward properties of this type of graphs:
Claim 48.
Let $G$, $\Delta\leq\Delta_{\max}$, $c$, and $\ell$ be as in Definition 47. Then for $\mathcal{B}(G)$ the following holds:
  • $\mathcal{B}(G)$ contains no parallel edges pointing in the same direction.
  • The maximum in-degree of $\mathcal{B}(G,c)$ is upper bounded by $2(\Delta_{max}+1)^{2}$.
  • The maximum out-degree of $v\in X_{2}$ is upper bounded by $2(\Delta_{max}+1)^{2}$.
Proof.
It is straight-forward to check that for any $u,v\in V(\mathcal{B}(G))$ at most one edge $(uv)$ can be inserted.
The second item follows from the fact that at most $(\Delta_{\max}+1)^{2}$ bichromatic paths can go through any vertex of $G$. More specifically, if $v\in X_{1}$ then at most $2$ in-edges can be incident to $v$, and if $v\in X_{2}$, then at most $(\Delta_{\max}+1)^{2}$ in-edge of $v$ can belong to $Y_{3}$ and at most $(\Delta_{\max}+1)$ in-edges of $v$ can belong to $Y_{2}$. In this case, no in-edges belong to $Y_{1}$.
Similarly, for the third item, we note that for $v\in X_{2}$ at most $(\Delta_{\max}+1)$ out-edges of $v$ belongs to $Y_{2}$ and at most $(\Delta_{\max}+1)^{2}$ out-edges of $v$ belongs to $Y_{1}$, since $v$ can be the endpoint of at most $(\Delta_{\max}+1)^{2}$ maximal bichromatic paths. No out-edges of $v$ belongs to $Y_{3}$. ∎
Finally, we note that any non-overlapping $i$-step process with step-length at most $\ell$ corresponds to a directed path in $\mathcal{B}(G)$ in a canonical way: simply map $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ to the path $(r(w_{1}),r(y_{1}),r(P_{1}),\dots,r(w_{i}),r(y_{i}),r(P_{i}))$.
Again let $G$, $\Delta\leq\Delta_{\max}$, $c$, and $\ell$ be as in Definition 47. We will maintain the data structures containing the following information:
  • We explicitly maintain the graph $\mathcal{B}(G)$.
  • For each edge $(x_{2},x_{1})\in Y_{1}$, we maintain for all $i\leq\sqrt{\log n}$ and for all neighbours $w$ of $y_{1}\in r^{-1}(x_{2})$, the following information in $\mathcal{D}_{i}(e,w)$: either all or the
    $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$
    shortest semi-consistent and augmenting $i$-step processes beginning at $w$ with $y_{1}\in r^{-1}(x_{2})$ and $P_{1}\in r^{-1}(x_{1})$.
In the next section, we will describe how to efficiently maintain the above information under colouring and un-colouring of edges. This is sufficient to easily construct a dynamic algorithm, since the above data structures readily contain information about which edges to recolour in order to extend the maintained colouring.
6.3 Maintaining the data structures
[…]
Maintaining the first item:
In order to explicitly maintain $\mathcal{B}(G)$, we will consider the $(\Delta_{\max}+1)^{2}$ graphs induced in $G$ by any pair of distinct colours $\kappa_{1},\kappa_{2}$. Since the colouring $c$ is at all times proper, these graphs consist only of isolated points, paths, and even cycles. In order to dynamically maintain the representation, we maintain each component in a Top-tree [AHdLT05]. In the case of even cycles, we cannot store the component directly in a dynamic tree data structure, but in that case we store the entire component except for one arbitrary edge in a Top-tree. As ancillary information, we store the number of degree $1$ vertices in the component. Note that there can be only 00, $1$, or $2$ vertices of degree $1$. We use this information to explicitly store $\mathcal{B}(G)$: whenever a maximal bichromatic path is changed, we remove the corresponding node and all its related edges. Next we insert representatives of any un-represented maximal bichromatic paths, and add the edges incident to these representatives as specified by Definition 47.
Maintaining the second item:
[…]
Next, we will cleanse the marked data structures and recompute their contents from scratch. We will re-compute the marked data structures using the invariant that all data structures of the form $\mathcal{D}_{i-1}$ should be recomputed before the first data structure of the form $\mathcal{D}_{i}$ is recomputed. Performing the re-computation in the case where $i=1$ is straight-forward as there only is one choice for the $1$-step process. To recompute the structures of the form $\mathcal{D}_{i}$ assuming that all structures of the form $\mathcal{D}_{j}$ with $j\leq i-1$ are correct, now amounts to the following: we perform the first step as prescribed in the description of the data structures. Next we use the data-structures of the form $\mathcal{D}_{i-1}$ along the bichromatic path to find the shortest suggestions of augmenting and consistent $(i-1)$-step processes (here the number of vertices we need to traverse to get to an extension is factored into the suggested length). Then we attempt each extension in increasing size of their lengths. For each attempted extension, we check that it is indeed consistent with our prefix extension by explicitly following the path along and checking the conditions as specified in Definition 15 and the following text. If the extension is not consistent, we don’t include it and attempt the next extension instead. We continue this process until we have enough extensions. It is crucial here that each $i-1$ structure contains enough suggested path that even if the maximum number of them overlap, the data structure can at least suggest the required number of non-overlapping paths. We show that this is indeed the case in the next section.
Performing $\texttt{colour}(e,\kappa)$ is now completely symmetric to the above. Assigning $e$ a colour $\kappa$ also causes some data structures to, potentially, be out-dated. These can now be fixed similarly to before. We provide a more in-depth description below:
Maintaining a bichromatic skeleton:
We will let $F_{\kappa_{1},\kappa_{2}}$ denote the pseudo-forest induced in $G$ by all edges coloured $\kappa_{1}$ or $\kappa_{2}$. For some vertex $v\in G$, we let $PT_{\kappa_{1},\kappa_{2}}(v)$ be the pseudo-tree in $F_{\kappa_{1},\kappa_{2}}$ containing $v$. We store every $PT_{\kappa_{1},\kappa_{2}}$ as a tree $T_{\kappa_{1},\kappa_{2}}$ stored as a Top-tree [AHdLT05] and possibly one extra edge stored in a special slot $S_{\kappa_{1},\kappa_{2}}$. We denote by $T_{\kappa_{1},\kappa_{2}}(v)$ the Top-tree currently containing $v$. Note that we will maintain the invariant that $S_{\kappa_{1},\kappa_{2}}=xy$ only if $x,y$ already belong to the same Top-tree i.e. if $T_{\kappa_{1},\kappa_{2}}(v)=T_{\kappa_{1},\kappa_{2}}(u)$. To perform BCS-colour$(uv,\kappa)$, we do as follows:
  • For all $\kappa^{\prime}\neq\kappa$, we check if $T_{\kappa_{1},\kappa_{2}}(v)=T_{\kappa_{1},\kappa_{2}}(u)$. If this is not the case, then we simply call $\texttt{link}(u,v)$ on the Top-tree data-structures. Otherwise, we store $uv$ in the special slot $S_{\kappa_{1},\kappa_{2}}$.
  • For each Top-tree we store as ancillary information the number of vertices with degree $1$ or 00 contained in each cluster, not counting the boundary vertices. It is straight-forward to maintain this information under join and split.
  • After each modification to the Top-tree, we perform a non-local search for each leaf and check if they are incident to the edge stored in $S_{\kappa_{1},\kappa_{2}}$ (if any such edge). If so, the vertices’ degrees are changed accordingly, and they might no longer be classified as vertices with degree $1$ or 00.
  • If the Top-tree has either $1$ or $2$ such vertices, then it is classified as a bichromatic path.
  • Next, we remove every vertex (and incident edges) representing a bichromatic path whose Top-tree, or related special slot, was modified in the above steps.
  • Finally, we add representatives (and incident edges) of all bichromatic paths whose Top-tree, or related special slot, was modified in the above steps.
  • To add the incident edges, we access the leaves of the Top-tree to find their neighbours in $G$ as well as the first $\ell$ vertices (from either direction) on the leaf-to-leaf path in the Top-tree. Finally, we add the edges as prescribed by Definition 47.
  • Finally, we add parallel, oppositely directed edges between $r(u)$ and $r(v)$.
To perform BCS-uncolour$(uv)$ we only have to modify the first step. The rest of the steps are exactly as before. Instead of calling $\texttt{link}(u,v)$, we call $\texttt{cut}(u,v)$ (or remove $uv$ from $S_{\kappa_{1},\kappa_{2}}$ if it stored there). Finally, if $S_{\kappa_{1},\kappa_{2}}=xy$ is non-empty, we call $\texttt{link}(x,y)$, and set $S_{\kappa_{1},\kappa_{2}}=\emptyset$. After this, we do the same clean-up step as we did in BCS-colour, except that insteado of adding parallel, oppositely directed edges between $r(u)$ and $r(v)$, we remove any such edges.
Maintaining the $\mathcal{D}_{i}$’s:
To perform un-colour$(uv)$ we do as follows:
  • First we un-colour $uv$ in $G$ and update $\mathcal{B}(G)$ as described earlier.
  • Next, we mark some of the edges of $\mathcal{B}(G)$ with the following scheme: an edge $e$ is marked $j$ if the data structures $\mathcal{D}_{i}(e,w)$ for $j\leq i\leq\sqrt{\log n}$ needs to be updated. An edge might be marked multiple times, but we need only keep the lowest mark, it receives.
  • All new edges in $\mathcal{B}(G)$ from step $1$ are marked $1$.
  • We let $R$ be the set containing representatives of the $1$-neighbourhood of $u$ and $v$ in $G$, all representatives of endpoints of new bichromatic paths of $\mathcal{B}(G)$, and all representatives of former endpoints of no longer-existing bichromatic paths.
  • All edges in $Y_{1}$ incident to $R$ are marked $1$.
  • We set $M_{1}=R$. We will mark all in-edges of vertices belonging to the set $N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{1}))$ also belonging to $Y_{1}$ i.e. the set $E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{1})))$ with $2$. Finally, we set $M_{2}=N_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{1})))$.
  • For all $1<i\leq\sqrt{\log n}$, we mark all in-edges of vertices belonging to the set $N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{i}))$ belonging to $Y_{1}$ with $i+1$. Finally, we set $M_{i+1}=N_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{i})))$.
  • Having updated all data-structures at level $<i$, we update $\mathcal{D}_{i}(e,w)$ as follows: To update $\mathcal{D}_{i}(x_{1}x_{2},w)$ for $i=1$, construct every augmenting $1$-step process starting at $w_{1}=r^{-1}(w)$ with $y_{1}=r^{-1(x_{2})}$ and $P_{1}\in r^{-1}(x_{1})$. If the process is augmenting, we save the vertices in the process. Note that it is crucial that this information can be efficiently computed without access to any other similar data structures.
    For $i>1$, we ensure that all of the data structures $\mathcal{D}_{j}(e,w)$ have been (re)computed for all $j<i$. To update $\mathcal{D}_{i}(x_{2}x_{1},w)$ we let $w_{1}=r^{-1}(w)$ and $y_{1}=r^{-1}(x_{2})$. We pick a valid choice of $P_{1}\in r^{-1}(x_{1})$ For each valid choice of $P_{1}$, there is a unique extension point $x=x(P_{1})$ corresponding to the last vertex of $P_{1}$ visited, if one traverses $P_{1}$ starting at $y_{1}$. Note that $x$ is among the first $\ell$ vertices of the maximal bichromatic path containing $P_{1}$. For each choice of $P_{1}$ (or equivalently $x$), we do as follows:
    1. 1.
      Construct a fan consistent with Lemma 13. We let $P_{2}(x)$ resp. $y_{2}(x)$ be the be bichromatic path resp. the endpoint of $P$ consistent with Lemma 13.
    2. 2.
      Next we consider $D_{i-1}(r(y_{2}(x))r(P_{2}(x),x)$ and extract the shortest consistent and augmenting $(i-1)$-step process stored here.
    3. 3.
      Then we check whether this extension is consistent with $P_{1}$ according to Definition 15. This is done by simply following the $(i-1)$-step process forward and checking the conditions. If not, then we are happy for now. Otherwise, we take the second smallest extension and repeat the process until we either exhaust all choices or we find a process and concatenate it with $r^{-1}(x)+y_{1}+P_{1}$ to form a consistent $i$-step process. We let $L_{x}$ denote the length of the above $i$-step process.
    4. 4.
      We repeat the above step for each choice of $x$, and then store the candidate chains in a BBST.
    5. 5.
      Then we extract the shortest chain, store its length and the vertices in the chain in the BBST representing $\mathcal{D}_{i}(x_{2}x_{1},w)$.
    6. 6.
      If the shortest chain went through $x$, then we repeat the above process to find the second shortest chain that does not overlap with $r^{-1}(x)+P_{1}(x)$ as before. Then $L(x)$ is defined in terms of this chain and again inserted into the BBST containing all of the shortest candidate chains.
    7. 7.
      This is repeated until $\mathcal{D}_{i}(x_{2}x_{1},w)$ contains the desired number of chains. If we at one point run out of chains to check, then we also terminate the above step and conclude that $\mathcal{D}_{i}(x_{2}x_{1},w)$ is updated.
    8. 8.
      Finally, we un-mark every marked edges.
In order to perform colour$(uv,\kappa)$, we do almost the same as above. The only change is that instead of un-colouring $uv$ in $G$, we colour it in $G$ with the colour $\kappa$. Then we update $\mathcal{B}(G)$ accordingly, and then proceed exactly as before.
6.4 Analysing the data structures
We start out by analysing the data structure for maintaining a bichromatic skeleton:
Lemma 49.
[…]
Proof.
[…]
After colouring $e$, the only new vertices of $X_{1}$ that need to be added are representatives of bichromatic paths containing $e$. Any bichromatic component $C$ containing $e$ can only contain edges with colours $\kappa$ and $\kappa^{\prime}$ for some $\kappa^{\prime}\neq\kappa$. Since the colouring maintained is forced to be proper, it follows that $C$ must be either an even cycle or a path (possibly of length 0). In particular, $C$ can always be represented by a pseudo-tree or a tree with one additional special edge. Therefore, the algorithm always correctly identifies the number of of vertices in $C$ of degree $1$ or 00. Hence, the algorithm will always identify if $C$ is a (possibly empty) bichromatic path, and correctly insert a representative of $C$ into $\mathcal{B}(G)$. So $X_{1}$ contains a representative of all current bichromatic paths. The algorithm also deletes any bichromatic path whose Top-tree representation was modified during the process above. Since $e$ is inserted into the representation of $G[\kappa,\kappa^{\prime}]$ for all $\kappa^{\prime}$, either the Top-tree or the related special slot representing the component of $e$ in $G[\kappa,\kappa^{\prime}]$ will be modified. In particular, either the Top-tree or the related special slot representing bichromatic paths no longer present will be modified and hence deleted. Therefore, $X_{1}$ is maintained correctly.
To see that the edge set is also correctly maintained, we note that the algorithm inserts the new edges in $Y_{2}$ as well as any relevant edges incident to new bichromatic paths. These are the only new edges that need to be added. It also removes any edges going to representatives no longer present.
To recover the update times, we note that any operation to the Top-trees can be performed in $O(\log n)$ time (see [AHdLT05]), and that any non-local search for edges or degree $1$ or 00 vertices can be done in $O(\log n)$ time [AHdLT05] per edge. It is also standard that the number of such vertices can be maintained with only $O(1)$ overhead. Since we update or search only $O(\Delta_{\max})$ Top-trees, the total time spend searching in the Top-trees is in $O(\Delta_{\max}\cdot{}\ell\cdot{}\log n)$. Furthermore, for each new (and old) bichromatic path, of which there are at most $O(\Delta_{\max})$, we have to add (and delete) $2\Delta_{\max}$ edges to (from) $Y_{1}$ and $2\ell$ edges to (from) $Y_{3}$. Each edge is easy to insert (or delete) in constant time after performing the searches in the Top-trees already accounted for. Hence this takes additional time at most $O(\Delta_{\max}^{2}+\Delta_{max}\cdot{}\ell\cdot{}\log n)$ and the lemma follows. ∎
Remark 50.
We remark that it is straight-forward to allow insertion and deletion of isolated vertices into $G$ in $O(1)$ update time per insertion/deletion.
Next we analyse the data structure for maintaining the $\mathcal{D}_{i}$’s. We will show the following Theorem:
Theorem 51.
Let $G$ be a dynamic graph, and let $\Delta_{\max}$ be a known upper-bound on the maximum degree throughout the entire update sequence. Then, we can maintain the data structures $\mathcal{D}_{i}(\cdot{},\cdot{})$ in $(\Delta_{\max}+1)^{\tilde{O}(\sqrt{\log n})}$ worst-case update time per operation.
We have already described how to update the data structure in Section 6.3. In the following, we show correctness and analyse the update times:
Correctness:
First we show that any data structure whose contents are no longer valid, will indeed be properly marked:
Lemma 52.
Assuming that all data-structures $\mathcal{D}_{i}$ are initially correct, then after the marking process of a call to $\emph{{colour}}(e,\kappa)$ or $\emph{{uncolour}}(e)$ concludes, we have that: if $f$ is any edge and $w$ any vertex with data-structure $\mathcal{D}_{i}(f,w)$ no longer correct, then $f$ is marked with $j\leq i$.
Proof.
Let $f=x_{2}x_{1}$, $w$ and $i$ be resp. an edge, a vertex and index such that $\mathcal{D}_{i}(f,w)$ is no longer correct. We will show that then $f$ is marked with $j\leq i$.
There are two reasons why $\mathcal{D}_{i}(f,w)$ might no longer be correct. The first reason $A)$ is that it contained an $i$-step process with step-length $\leq\ell$ which used a changed bichromatic path and so is no longer valid. Similarly it could also be that a new bichromatic path allows for a shorter consistent and augmenting $i$-step process with step-length $\leq\ell$, and this needs to be reflected in $\mathcal{D}_{i}(f,w)$. The second reason $B)$ is that changes to the colouring $c$ causes some invocations of Lemma 13 to be altered and so some of the stored processes are no longer consistent. Of course both reasons could be present at the same time. Observe that it is sufficient to show that if any $i$-step process with step-length $\leq\ell$ beginning with $w+x_{2}+x_{1}$ can reach either 1) a changed bichromatic path in $(i-1)$-steps or less or 2)2)2 ) any vertex in the one-hop neighbourhood of $e$ in $(i-1)$-steps or less, then $f$ is marked with $j\leq i$. Indeed, if 1)1)1 ) does not hold, then the reason $A)$ can never occur. Similarly, if 2)2)2 ) does not hold, then the reason $B)$ cannot occur.
Thus it is sufficient to show that if 1)1)1 ) or 2)2)2 ) holds then $f\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$ for some $2\leq j\leq i$ or $f$ is incident to a vertex in $R$. Indeed, then $f$ is marked with $j\leq i$.
We begin by assuming that 1)1)1 ) holds. This means that after the changes to $c$, there is an $i$-step process $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ with $w_{1}=w$, $y_{1}=r^{-1}(x_{2})$, and $P_{1}\in r^{-1}(x_{1})$ using a changed bichromatic path $P$ during the $j^{\text{th}}$ step i.e. $P_{j}$ is a sub-path of $P$. Let $j\leq i$ be the smallest number for which the above is true.
This means that before the change to $c$, $w_{1}+y_{1}+P_{1}+\dots+w_{j}+y_{j}$ was also an initial segment of some $j$-step process. Hence, this initial segment corresponds to a directed path in $\mathcal{B}(G)$ of the form $(r(w_{1}),r(y_{1}),r(P_{1}),\dots,r(w_{j}),r(y_{j})$ present both before and after the update to $c$. This follows from the minimality of $j$ and the discussion directly after Claim 48. Observe that then $y_{j}\in R$, $(r(w_{j})r(y_{j}))\in Y_{2}$, $(r(P_{j-1})r(w_{j}))\in Y_{3}$. Hence, if $j=1$, then $f=(x_{2}x_{1})=(r(y_{1})r(P_{1}))$ is incident to $R$, and so we would be done. More generally note that for all $k\leq j-1$, we have that $(r(y_{k})r(P_{k}))\in Y_{1}$, $(r(w_{k})r(y_{k}))\in Y_{2}$, and if $k>1$ that $(r(P_{k-1})r(w_{k}))\in Y_{3}$. Thus if $\psi:\mathcal{P}(V(\mathcal{B}(G)))\rightarrow\mathcal{P}(V(\mathcal{B}(G)))$ maps a set $A\in\mathcal{P}(V(\mathcal{B}(G)))$ to the set $N_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(A)))$, then $y_{k-1}\in\psi(\{y_{k}\})$ for $1<k\leq j$. In particular, it follows by induction that $y_{1}\in\psi^{j-1}(\{y_{j}\})$. So if $y_{j}\in R$, then $y_{1}\in M_{j}$ and therefore $f=(r(y_{1})r(P_{1}))\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$.
On the other hand, assume that 2)2)2 ) holds. Then there is an $i$-step process $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ with $w_{1}=w$, $y_{1}=r^{-1}(x_{2})$, and $P_{1}\in r^{-1}(x_{1})$ such that for some $j$ we have $w_{j}=r(z)$ for some $z\in N^{1}(e)\cup e$ in the coloured inclusive neighbourhood of $z$. Indeed, otherwise the application of Lemma 13 cannot be changed at $z$, as it depends only on the colours of edges incident to vertices in the $1$-hop neighbourhood of $z$. Therefore there exist a vertex $z^{\prime}$ in the coloured neighbourhood of $z$ whose representative $r(z^{\prime})$ belongs to $R$. WLOG we can assume that $j$ is the smallest such index among all choices.
If $j=1$, then we are now done as $r^{-1}(y_{1})$ must be one such vertex and therefore $f$ is incident to a vertex in $R$. For $j>1$ we can conclude, similarly to case 1)1)1 ), that $y_{1}\in\psi^{j-1}(R)$ and therfore that $f\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$ and thusly marked with $j$. This concludes the proof of the lemma. ∎
Next we show that assuming that all un-marked data structures are maintained correctly, then we will update all marked structures correctly. Together with Lemma 52 above, this shows correctness of the update-algorithm.
Lemma 53.
[…]
Proof.
[…]
The base case is simple, as we just check all possible choices for semi-consistent and augmenting $1$-step processes with step-length at most $\ell$. At most one such process can exist, and if so, we store it. Note that this information can be computed without requiring information stored at any other data structures of the form $D_{\cdot}(\cdot,\cdot)$.
For the induction step, we assume that all data-structures $D_{k}(f,z)$ are correct for $k<i$. We fix choices of $f^{\prime}$ and $z^{\prime}$, and show that $D_{i}(f^{\prime},z^{\prime})$ is correctly computed. We first observe that the values $L(x)$ correctly lower bound the shortest semi-consistent and augmenting $i$-step process with step-length at most $\ell$ extended through $x$ at the $2^{\text{nd}}$ step.
Claim 54.
Let $f^{\prime}=(x_{2}x_{1}),z^{\prime}$ be as above. Let $x$ be any extension-point corresponding to some $P_{1}\in r^{-1}(x_{1})$. Then, at all times $L(x)$ is a lower-bound for the shortest semi-consistent and augmenting $i$-step process with step-length at most $\ell$ beginning with $z^{\prime}+r^{-1}(x_{2})+P_{1}+x$ not already present in $D_{i}(f^{\prime},z^{\prime})$.
Proof.
Any semi-consistent and augmenting $i$-step process with step-length at most $\ell$ beginning with $z^{\prime}+x_{2}+P_{1}+x$ must be extended by a semi-consistent and augmenting $(i-1)$-step process with step-length at most $\ell$ beginning with $x+y_{2}+P_{2}$ for some sub-path $P_{2}$ of $P$, where $y_{2}$ and $P$ are chosen consistently with Lemma 13, that furthermore does not overlap $z^{\prime}+x_{2}+P_{1}$. This is straight-forward to verify via Definition 15. By the induction hypothesis, we have that whenever $L(x)$ is set, it is set to $|P_{1}|+1+$ the length of the shortest semi-consistent and augmenting $(i-1)$-step process $C$ with step-length at most $\ell$ that satisfies the following:
  1. 1.
    $C$ begins with $x+y_{2}+P_{2}$ for some sub-path $P_{2}$ of $P$, where $y_{2}$ and $P$ are chosen consistently with Lemma 13.
  2. 2.
    $C$ is shortest among such processes that is not either already represented in $D_{i}(f^{\prime},z^{\prime})$ or has already been confirmed to overlap $z^{\prime}+x_{2}+P_{1}$ in the sense of Definition 15.
In particular, it is clear that $L(x)$ therefore lower bounds the length of any sought-after process. ∎
Secondly, we observe that if we have $\alpha$ semi-consistent and augmenting $(i-1)$-step processes with step-length at most $\ell$ beginning with $x+y_{2}+P_{2}$ for some $P_{2}$ with the choices of $y_{2}$ and $P_{2}$ consistent with Lemma 13, then $D_{i-1}((r(y_{2})r(P_{2}),r(x))$ contains at least $\min\{\alpha,\sqrt{\log n}(\sqrt{\log n}-(i-1))\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell\}$ augmenting and semi-consistent $(i-1)$-step processes which can be added to the $1$-step process $z+r^{-1}(x_{2})+P_{1}$ to form either all or the shortest $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ semi-consistent and augmenting processes of the form we seek.
Indeed, we note that by applying Lemma 41 to $z^{\prime}+r^{-1}(x_{2})+P_{1}$, we find that at most $i\cdot{}3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}\ell$ of all non-overlapping $j$-step processes with $j\leq i-1$ can overlap with $z^{\prime}+r^{-1}(x_{2})+P_{1}$ in the sense of Definition 15. Note that this bound also holds for the extensions stored in $D_{i-1}((r(y_{2})r(P_{2}),r(x))$, as the first path of any extension can only overlap the $2$-hop neighbourhood of $z^{\prime}$, in which case the extension is accounted for above via Lemma 41, or via a later part of the process that loops back and overlaps the two hop-neighbourhood of $z^{\prime}+r^{-1}(x_{2})+P_{1}$. But this later part of the process then forms a $j$-step process for $j<i-1$ that starts outside of the two hop-neighbourhood of $z^{\prime}+r^{-1}(x_{2})+P_{1}$ and then reaches the two hop-neighbourhood of $z^{\prime}+r^{-1}(x_{2})+P_{1}$. Thus it is also accounted for by Lemma 41.
In particular, since $D_{i-1}((r(y_{2})r(P_{2}),r(x))$ contains the shortest
$\sqrt{\log n}(\sqrt{\log n}-(i-1))\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$
candidate extensions (or fewer, if fewer exists), we either find all extensions that will not overlap with $z^{\prime}+r^{-1}(x_{2})+P_{1}$ in the sense of Definition 15 or we find $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ such extensions. In particular, the above together with Claim 54 implies that the algorithm will locate either all or the $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ shortest semi-consistent and augmenting $i$-step processes with step-length at most $\ell$ beginning with $z^{\prime}+r^{-1}(x_{2})+P^{\prime}$, where $P^{\prime}\in r^{-1}(x_{1})$. This concludes the induction step. ∎
Lemma 53, Lemma 52, and Lemma 49 together show correctness of the data structure and the update algorithms for the supported operations.
Implementation and analysis:
We first briefly comment on implementations certain data structures. For each BBST, we can use Red-blacks trees [CLRS22] to support basic operations in $O(\log n)$ worst-case time. For Top-trees we use the version from [AHdLT05]. As for the analysis, we first show that not too many edges are marked:
Lemma 55.
In each invocation of the marking scheme, at most $2^{3i}(\Delta_{\max}+1)^{6i}|R|$ edges are marked with $i$. Hence $|M_{i}|\leq 2^{3i}(\Delta_{\max}+1)^{6i}|R|$.
Proof.
The proof is by induction on $i$. For $i=1$, we note that by item two and item three of Claim 48 at most $4(\Delta_{\max}+1)^{2}\cdot{}|R|$ edges are initially marked $1$. Also $|M_{1}|=|R|$. For the induction step, we assume the statement for all $j\leq i-1$. We note that applying item two of Claim 48 thrice shows that $M_{i}\leq 2^{3}(\Delta_{\max}+1)^{6}|M_{i-1}|$ and that the number of edges marked with $i$ is also at most $2^{2}(\Delta_{\max}+1)^{6}|M_{i-1}|$. Applying the induction hypothesis now yields the result. ∎
Next we show that $R$ is not too big:
Lemma 56.
In each invocation of the marking scheme, we have that $|R|\leq 2+2(\Delta_{\max}+1)+6(\Delta_{\max}^{2}+1)^{2}\leq 10(\Delta_{\max}^{2}+1)^{2}$.
Proof.
We note that at most $3(\Delta_{\max}^{2}+1)^{2}$ bichromatic paths are altered after a removal or insertion of a coloured edge: namely 3 for each pair of colours; the bichromatic path that used to contain the edge and the two new bichromatic paths that this path was split into (for insertion it is of course opposite; one new bichromatic path that is the fusion of two old). There are at most $(\Delta_{\max}+1)^{2}$ choices for such colour pairs. Note that each path has at most $2$ endpoints. Finally, there are at most $2(\Delta_{\max}+1)$ vertices in the one-hop neighbourhood of $u$ and $v$. Adding these quantities gives the bound. ∎
These two lemmas allow us to analyse the time it takes to mark all edges:
Lemma 57.
It takes $\tilde{O}(2^{3(\sqrt{\log n}+4)}(\Delta_{\max}+1)^{6(\sqrt{\log n}+2)})$ time to mark all edges.
Proof.
We can mark each edge with constant overhead by performing a backwards BFS search along the relevant edges (assuming we store different types of edges in different BBST’s). Hence, the Lemma follows by combining Lemma 55 with Lemma 56, and summing over all choices of $1\leq i\leq\sqrt{\log n}$:
$10(\Delta_{\max}+1)^{2}\sum\limits_{i=1}^{\sqrt{\log n}}2^{3i}(\Delta_{\max}+1)^{6i}$ $\leq 10(\Delta_{\max}+1)^{2}\cdot{}2^{3(\sqrt{\log n}+1)}(\Delta_{\max}+1)^{6(\sqrt{\log n+1})}$
$\leq 2^{3(\sqrt{\log n}+4)}(\Delta_{\max}+1)^{6(\sqrt{\log n}+2)}$
Next, we bound the time it takes to check whether an extension is overlapping:
Lemma 58.
Given a $1$-step process $w_{1}+y_{1}+P_{1}$, with $r^{-1}(w_{2})$ the second last vertex of $P_{1}$, and a semi-consistent and augmenting $(i-1)$-step process $w_{2}+y_{2}+P_{2}+\dots+w_{i}+y_{i}+P_{i}$, we can check if $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ is a semi-consistent and augmenting $i$-step process in $O((i+1)\cdot{}\ell\cdot{}(\Delta_{\max}+1)^{3})$ time.
Proof.
We check whether $w_{2},y_{2}$ and $P_{2}$ is consistent with Lemma 13 in $\tilde{O}(\Delta_{\max}^{3})$. Next we can can trace and temporarily mark the two-hop neighbourhood of every vertex in $w_{1}+y_{1}+P_{1}$, as well as the edges in $P_{1}$. Finally, we can follow along the $(i-1)$-step process and check the conditions of Definition 15 by using the temporary marks. It takes $\tilde{O}((i+1)\ell)$ time to trace the process and $\tilde{O}(\ell(\Delta_{\max}+1)^{2})$ to temporarily mark all of the edges and vertices. We note here that any update time required to correctly locate edges or vertices in memory is captured by the $\tilde{O}$-notation which accounts for look-ups in BBST. ∎
We have already seen in the proof of Lemma 53 that not too many of the processes stored in neighbouring data structures can be overlapping. This is captured in the following Lemma:
Lemma 59.
Suppose we are given a $1$-step process $w_{1}+y_{1}+P_{1}$, with $r^{-1}(w_{2})$ the second last vertex of $P_{1}$. Let $C^{\prime}$ be a semi-consistent and augmenting $(i-1)$-step processes of the form $C^{\prime}=w_{2}+y_{2}+P_{2}+\dots+w_{i}+y_{i}+P_{i}$. Suppose further that the choice of $w_{2},y_{2}$ and $P_{2}$ is consistent with Lemma 13. Then at most $i\cdot{}3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}\ell$ choices of $C^{\prime}$ can cause $C=w_{1}+y_{2}+P_{2}+\dots+w_{i}+y_{i}+P_{i}$ to not be a semi-consistent and augmenting $i$-step process.
Proof.
Since the extension is already assumed to be consistent with Lemma 13, we need only check that $C^{\prime}$ does not overlap $w_{1}+y_{1}+P_{1}$ in the sense of Definition 15. We can however bound the number of such $C^{\prime}$ reaching the two-hop neighbourhood of $w_{1}$ by applying Corollary 42 to every point in this neighbourhood. We note that due to the extension being consistent with Lemma 13, any other violations has to be because $P_{j}$ intersects the $2$-hop neighbourhood of $w_{1}+y_{1}+P_{1}$ for some $j\geq 3$ or because we reuse the maximal bichromatic path containing $P_{1}$. In the first-case the non-overlapping process $C^{\prime}=w_{3}+y_{3}+P_{3}+\dots+w_{i}+y_{i}+P_{i}$ also intersects the $2$-hop neighbourhood of $w_{1}+y_{1}+P_{1}$. But since $w_{3}$ is disjoint from the $2$-hop neighbourhood of $w_{1}+y_{1}+P_{1}$, this process is accounted for by Corollary 42. The second case is similarly accounted for by Corollary 42. ∎
Finally, we analyse the time it takes to update the data structures:
Lemma 60.
Let $f$ be marked with $j\leq i$. Then for any $w$, we can update $\mathcal{D}_{i}(f,w)$ in $\tilde{O}(3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+4}\cdot{}\ell^{3}\cdot{}\log^{2}{n})$ time.
Proof.
Note firstly that the $\tilde{O}$-notation takes care of any $\log n$ factors arising from inserting, searching or deleting things from BBST’s. Here it is important to note that the BBST’s used to update the data structures never have sizes greater than $\tilde{O}(\ell^{2}\cdot{}\log{n}\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2})$. Indeed, that is the maximum number of chain extensions consistent with Lemma 13, stored in the neighbourhoods of the first $\ell$ vertices along $P_{1}$.
Secondly, Lemma 59 implies that at most $\sqrt{\log n}\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ of these extensions can be overlapping with the first-step of the process in the sense of Definition 15.
Hence, at most $3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell\cdot{}\log n$ such chains need to be checked, before $\mathcal{D}_{i}(f,w)$ is either full or contains all semi-consistent and augmenting $i$-step processes. Each extension can be checked in $\tilde{O}((i+1)\cdot{}\ell\cdot{}(\Delta_{\max}+1)^{3})$ time by Lemma 58 and since it only takes $\tilde{O}(\ell\Delta_{\max}^{3})$ to compute and locate the correct choices of $\mathcal{D}_{i-1}(\cdot{},\cdot{})$ to look for extension suggestions in. By noting that $i\leq\sqrt{\log n}$, the Lemma follows. ∎
Combining all of the above finalises the proof of Theorem 51.
6.5 The proof of Theorem 5
In this section, we present a proof of Theorem 5 (restated below for convenience).
Theorem 61 (Identical to Theorem 5).
Let $G$ be a dynamic graph, and let $\Delta_{\max}$ be a known upper-bound on the maximum degree throughout the entire update sequence. Then, we can maintain a proper $(\Delta_{\max}+1)$-edge-colouring of $G$ in $(\Delta_{\max}+1)^{\tilde{O}(\sqrt{\log n})}$ worst-case update time per operation.
Description of the algorithm:
[…]
  • Call $\texttt{add}(e)$.
  • Then we pick an arbitrary vertex $u\in e$, and construct a consistent fan according to Lemma 12. The fan ends at a vertex $y_{1}$ and outputs a bichromatic path $P$.
  • If $|P|\leq\ell$, we have discovered an augmenting chain, and we can immediately jump to step 5)5)5 ). Otherwise, we look in $\mathcal{D}_{i}(r(y_{1})r(P),u)$ for all $i$ in order to find a consistent and augmenting $j$-step process $C$ for some $j$. Note that Corollary 46 and Theorem 51 guarantees that such a chain exists.
  • According to Lemma 18, $C$ guarantees some augmenting and consistent $j$-step Vizing chain. We can explicitly compute this chain by tracing the process and computing the consistent fans according to Lemma 13.
[…]
  • Finally since the chain is augmenting, after shifting it the remaining un-coloured edge $f$ has a free colour $\kappa$. We colour it using $\texttt{colour}(f,\kappa)$.
Correctness:
[…]
Analysis:
[…]
7 Putting it all together
In this Section, we combine the results of the previous section with a degree-scheduler from [DHZ19] to show Theorem 2
Theorem 62 (Identical to Theorem 2).
Let $G$ be a dynamic graph, and $10<\varepsilon<$ 1 a given parameter. We can maintain a proper $((1+\varepsilon)\Delta)$-edge-colouring of $G$ in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ amortised update time per operation.
7.1 The final algorithm
For now, we will still assume that we have access to some upper bound $\Delta_{\max}$ on the maximum degree of $G$ throughout the entire update sequence. We will then get rid of this assumption in the next section by using a degree-scheduler due to Duan, He, and Zhang [DHZ19].
The algorithm is now the following: If $\Delta_{\max}\leq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$: we simply apply Theorem 5 directly. Since $\Delta_{\max}+1\leq\lceil(1+\varepsilon)\Delta_{\max}\rceil$ for any $\varepsilon>0$, we are immediately done by Theorem 5.
Otherwise, if $\Delta_{\max}>\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$, we first apply Theorem 4 to get a degree-splitting hierarchy of height $0<h\leq 2\sqrt{\log n}$. Then we apply Theorem 5 with the upper bound on the maximum degree equal to $\Delta^{\prime}_{\max}$, where $\Delta^{\prime}_{\max}=\hat{\Delta}_{h}$, to each graph of $\mathcal{G}_{h}$, using a distinct palette to colour each graph. Correctness of the algorithm then follows immediately from Theorem 4 and Theorem 5.
As for the update time of the algorithm, we note that by Theorem 4 each sequence of $\sigma$ updates to $G$ results in at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ updates to graphs in $\mathcal{G}_{h}$. We can handle each such update in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time by Theorem 5, so in total, we process the $\sigma$ updates in $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time, thus resulting in an amortised update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time, as claimed.
Finally, we analyse the number of colours used. We note that by Theorem 4, there exists non-negative integers $i,j$ with $i+j=h$ such that:
$|\mathcal{G}_{h}|\leq t_{1}^{i}t_{2}^{j}$
and
$\Delta^{\prime}_{\max}=\hat{\Delta}_{h}\leq(1+\tfrac{\varepsilon}{16})\frac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
In particular, each graph in $\mathcal{G}_{h}$ is coloured using at most $(1+\tfrac{\varepsilon}{16})\frac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}+1$ colours. Therefore, the total number of colours used is upper bounded by:
$|\mathcal{G}_{h}|\cdot{}((1+\tfrac{\varepsilon}{16})\frac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}})+1)\leq(1+\tfrac{\varepsilon}{2})\Delta_{max}+t_{1}^{i}t_{2}^{j}\leq(1+\varepsilon)\Delta_{\max}$
Here we used the fact that Theorem 4 terminates with a hierarchy for which $\Delta_{\max}\geq\tfrac{2\cdot{}t_{1}^{i}t_{2}^{j}}{\varepsilon}$.
7.2 Dynamic $\Delta$
In this subsection, we finish the proof of Theorem 2. This last part contains no new ideas, as we can use the degree-scheduler of Duan, He and Chang [DHZ19] essentially ad verbatim. For the sake of completeness, we briefly sketch their approach below.
Let $\mu$ be some constant not specified until later (we will ultimately set $\mu=\Theta(\varepsilon)$). Consider the exponentially increasing intervals $I_{i}:=[(1+\mu)^{i-1},(1+\mu)^{i}]$ for $i\in\{0,1,\dots,s\}$ with $s$ minimum subject to $(1+\mu)^{s}\geq n$. For each $i$, we will maintain a subgraph of $G$, say $G_{i}$ with the guarantees that if $\Delta(G)\in I_{i}$, then $G_{i}=G$, and that for all $i$, we have $\Delta(G_{i})\leq(1+\mu)^{i+1}$ always. Furthermore, we will at all times maintain some index $j$ for which $\Delta(G)\in I_{j}$.
Given the above, one can run any $(1+\varepsilon^{\prime})\Delta_{\max}$-edge-colouring algorithm on all of the $G_{i}$’s, and then let the colouring of $G_{j}$ be the currently valid edge-colouring. If one sets $\Delta_{\max,i}=(1+\mu)^{i+1}$ in the instance of the edge-colouring algorithm on $G_{i}$, one has the guarantee that at all times the algorithm maintains a $(1+\mu)^{2}(1+\varepsilon^{\prime})\Delta$-edge-colouring. Indeed, at all times in the copy currently indexed by $j$ will have $\Delta\cdot{}(1+\mu)^{2}\geq\Delta_{\max,j}$. Setting $\varepsilon^{\prime},\mu\leq\tfrac{\varepsilon}{3}$ is then sufficient to achieve the correct number of colours.
In order to maintain this information, Duan, He and Chang do as follows: let $\delta(e)$ be the maximum degree of an endpoint of $e$. If $\delta(e)\leq(1+\mu)^{i}$, we add $e$ to $G_{i}$. Now $e$ stays in $G_{i}$ until $\delta(e)>(1+\mu)^{i+1}$, at which point in time it is removed, and not added back until $\delta(e)\leq(1+\mu)^{i}$ again. It is now easy to check that that every time some vertex $w$ crosses the threshold $(1+\mu)^{k}$ for some $k$, it performs at most $deg(w)$ insertions into $G_{k}$ or at most $deg(w)$ deletions from $G_{k}$. Between any set of insertions into some $G_{k}$, $deg(w)$ must have increased beyond $(1+\mu)^{k+1}$ and back down to $(1+\mu)^{k}$. All of the deletions incident to $w$ that decreased the degree from $(1+\mu)^{k+1}$ to $(1+\mu)^{k}$ can be charged the cost of performing the $deg(w)$ insertions into $G_{k}$. Deletions can be analysed in a symmetric way. In total, the amortised recourse of performing these changes is upper bounded by $O(\mu^{-1})$ for each update. Since there are at most $O(\tfrac{\log n}{\mu})$ choices of $i$, the total amortised recourse and update time becomes $O(\tfrac{\log n}{\mu})$. One can easily keep track of $j$ or directly calculate $j$ using knowledge of the current maximum degree $\Delta$. Therefore picking $\mu,\varepsilon^{\prime}=\theta(\varepsilon)$ with $\mu,\varepsilon^{\prime}\leq\tfrac{\varepsilon}{3}$ yields the corollary.
8 Acknowledgements
The author would like to thank Eva Rotenberg for the support and encouragements.